### v2

# Limits of open ASEP stationary measures near a boundary

Abstract
Consider the stationary measure of open asymmetric simple exclusion process (ASEP) on the lattice $\{1,\dots,n\}$. Taking $n$ to infinity while fixing the jump rates, this measure converges to a measure on the semi-infinite lattice. In the high and low density phases, we characterize the limiting measure and provide bounds on the convergence rates in total variation distance. Our approach involves bounding the total variation distance using generating functions, which are further estimated through a subtle analysis of the atom masses of Askey–Wilson signed measures.
82C22,
60K35,
82B23,
Open ASEP,
Askey–Wilson signed measures,
keywords:
[class=MSC]
keywords:
[…]
$0$-0003-2551-7965
1 Introduction
1.1 Preface
The open asymmetric simple exclusion process (ASEP) serves as a fundamental model for nonequilibrium systems with open boundaries and for Kardar–Parisi–Zhang (KPZ) universality. Over the past five decades, extensive studies have been dedicated to understand the stationary measure of open ASEP, encompassing a wide range of its asymptotic and limiting behaviors, including the particle densities [14, 28, 24, 20, 27, 15, 16, 29, 7, 30], limit fluctuations [13, 7, 30], large deviations [15, 16, 9] and open KPZ limits [12, 6, 3, 5]. See survey papers [4, 11] and more references therein. A significant portion of these studies stems from the matrix product ansatz (MPA) method introduced in the seminal work [14]. This method is notably related to the Askey–Wilson polynomials [29, 10] and processes [9].
We are interested in a straightforward limit of the open ASEP stationary measures near the boundary. Specifically, we will consider the stationary measure on the lattice $\{1,\dots,n\}$. We fix all the parameters $q,\alpha,\beta,\gamma,\delta$ of the model and take the system size $n$ to infinity. It is known from [23, 31] that such sequences weakly converge, and that the limiting probability measures on $\{0,1\}^{\mathbb{Z}_{+}}$ are stationary measures of certain ASEP systems on the semi-infinite lattice $\mathbb{Z}_{+}$ with parameters $q,\alpha,\gamma$. We mention that the stationary measures of this semi-infinite ASEP are not unique and are parameterized by the limiting densities at infinity. The aforementioned limits from finite lattices $\{1,\dots,n\}$ to $\mathbb{Z}_{+}$ were first studied by Liggett [23] assuming the so-called Liggett’s condition, see for example [23, Theorem 1.8 and Theorem 3.10]. Later in Grosskinsky [21, Theorem 3.2] and Sasamoto–Williams [31], a matrix product ansatz was developed to characterize the limiting measures, enabling the studies of their large deviations in Duhart [17] and Duhart–Mörters–Zimmer [18]. The limiting measures were further characterized in Bryc–Wesołowski [9, Theorem 12] in terms of the Askey–Wilson processes, in the ‘fan region’ part of the phase diagram.
In this paper we achieve two main goals. Firstly, using the Askey–Wilson signed measures introduced in a recent work Wang–Wesołowski–Yang [30], we characterize the limiting probability measures on $\{0,1\}^{\mathbb{Z}_{+}}$ in the ‘shock region’ part of the phase diagram. This complements the characterization Bryc–Wesołowski [9, Theorem 12] in the fan region and could serve as a useful tool for further studies of the asymptotics. Secondly, we investigate the following natural question:
Question 1.1.
At which scale of the (leftmost) sublattice does this convergence occur in total variation distance? To be specific, when comparing the open ASEP stationary measure on $\{0,1\}^{n}$ with the limiting measure on $\{0,1\}^{\mathbb{Z}_{+}}$ by measuring the total variation distance of their marginal distributions on the sublattice $\{1,\dots,m_{n}\}$, under which growth rate of the sequence $m_{n}$, $n=1,2,\dots$ does this total variation distance converge to zero?
In the case $\gamma=\delta=0$ and within the low density phase of the shock region, a recent work by Nestoridi and Schmid [25, Theorem 1.4] provides a growth rate. In this paper, we contribute another (partial) answer to this question: In both the high and low-density phases, the convergence occurs in total variation distance on the leftmost sublattice with a scale of $n/\log n$. We note that in the high density phase, the limiting measures on $\{0,1\}^{\mathbb{Z}_{+}}$ are in general no longer product Bernoulli measures (as in the low density phase), necessitating a different method.
Our approach involves bounding the total variation distance between two probability measures on $\{0,1\}^{m}$ by the values of their joint generating functions at specific points. To bound the generating functions, subtle estimations on the total variations of certain Askey–Wilson signed measures are necessary, which are derived through careful analysis of the masses of all the atoms.
1.2 Model and results
The open ASEP is a continuous-time irreducible Markov process on the state space $\{0,1\}^{n}$ with parameters
$\alpha,\beta>0,\quad\gamma,\delta\geq 0,\quad 0\leq q<1,$ (1.1)
which models the evolution of particles on the lattice $\left\{1,\dots,n\right\}$. In the bulk of the system, particles move at random to the left with rate $q$ and to the right with rate $1$. At the left boundary, particles enter at random with rate $\alpha$ and exit at random with rate $\gamma$. At the right boundary, particles enter at random with rate $\delta$ and exit at random with rate $\beta$. Any move of a particle is prohibited if the target site is already occupied. See Figure 1 for an illustration.
[…]
Figure 1: Jump rates in the open ASEP.
[…]
Figure 2: Phase diagrams for the open ASEP stationary measures. LD, HD, MC respectively stand for the low density, high density and maximal current phases.
We will work with a re-parameterization of the open ASEP system by $A,B,C,D$ and $q$, where
$A=\phi_{+}(\beta,\delta),\quad B=\phi_{-}(\beta,\delta),\quad C=\phi_{+}(\alpha,\gamma),\quad D=\phi_{-}(\alpha,\gamma),$ (1.2)
and
$\phi_{\pm}(x,y)=\frac{1}{2x}\left(1-q-x+y\pm\sqrt{(1-q-x+y)^{2}+4xy}\right),\quad\mbox{for }\;x>0\mbox{ and }y\geq 0.$ (1.3)
The quantities $\frac{A}{1+A}$ and $\frac{1}{1+C}$ have nice physical interpretations as the ‘effective densities’ near the left and right boundaries of the system, see for example a survey [11, Section 6.2].
One can check that (1.2) gives a bijection between (1.1) and
$A,C\geq 0,\quad-1<B,D\leq 0,\quad 0\leq q<1.$ (1.4)
We will assume (1.1) and consequently, (1.4) throughout the paper.
It is known since [14] that as the system size $n\rightarrow\infty$, the asymptotic behavior of open ASEP is governed by parameters $A$ and $C$, which exhibits a phase diagram (Figure 2) involving three phases:
  1. $\bullet$
    (maximal current phase) $A<1$, $C<1$,
  2. $\bullet$
    (high density phase) $A>1$, $A>C$,
  3. $\bullet$
    (low density phase) $C>1$, $C>A$.
There are also two regions on the phase diagram distinguished by [15, 16]:
  1. $\bullet$
    (fan region) $AC<1$,
  2. $\bullet$
    (shock region) $AC>1$.
We denote by $\mu_{n}:=\mu_{n}^{(A,B,C,D)}$ the (unique) stationary measure of open ASEP, which is a probability measure on $(\tau_{1},\dots,\tau_{n})\in\left\{0,1\right\}^{n}$, where $\tau_{i}\in\left\{0,1\right\}$ is the occupation variable on site $i$, for $i=1,\dots,n$. For any $1\leq m\leq n$, the marginal distribution of $\mu_{n}$ on the sublattice $\{1,\dots,m\}$ is denoted by $\mu_{n|m}:=\mu_{n|m}^{(A,B,C,D)}$, which is a probability measure on $(\tau_{1},\dots,\tau_{m})\in\{0,1\}^{m}$. For any two probability measures $\kappa$ and $\kappa^{\prime}$ on $\{0,1\}^{m}$, we denote their total variation distance by
$\mathsf{d}_{TV}(\kappa,\kappa^{\prime}):=\frac{1}{2}\sum_{x\in\{0,1\}^{m}}|\kappa(x)-\kappa^{\prime}(x)|=\max_{A\subseteq\{0,1\}^{m}}|\kappa(A)-\kappa^{\prime}(A)|.$
We equip $\{0,1\}^{\mathbb{Z}_{+}}$ with the infinite product $\sigma$-algebra. The measure $\mu_{n}$ on $\{0,1\}^{n}$ can be regarded as a probability measure on $\{0,1\}^{\mathbb{Z}_{+}}$ by setting $\tau_{i}=0$ for $i\geq n+1$, which we also denote by $\mu_{n}$. As mentioned in the preface, we will characterize the weak limit $\mu_{\infty}$ of measures $\mu_{n}$ as $n\rightarrow\infty$. We will also provide a bound of total variation distance between probability measures $\mu_{n|m}$ and the marginal of $\mu_{\infty}$ on the sublattice $\{0,1\}^{m}$, for $1\leq m\leq n$. This total variation bound will in particular imply that for sequences $m_{n}$, $n=1,2,\dots$ of growth rate $n/\log n$, the total variation distance tends to 00.
We now state our main theorem in the low density phase:
Theorem 1.2.
In the low density phase $C>A$, $C>1$, as $n\rightarrow\infty$, the open ASEP stationary measure $\mu_{n}$ weakly converges to the product Bernoulli measure on the semi-infinite lattice $\mathbb{Z}_{+}$ with density $\frac{1}{1+C}$.
We furthermore assume that $A/C\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $A\geq 1$. Then there exists $H>0$ depending on $A,B,C,D$ and $q$ and $\theta\in(0,1)$ depending on $A,C$ and $q$ such that for any $1\leq m\leq n$ we have:
$\mathsf{d}_{TV}\left(\mu_{n|m},\operatorname{Ber}_{m}\left(\frac{1}{1+C}\right)\right)\leq\theta^{n}(Hm)^{3m},$ (1.5)
where we use $\operatorname{Ber}_{k}(\rho)$ to denote the product Bernoulli measure on the lattice $\{1,\dots,k\}$ with density $\rho$, for $\rho\in[0,1]$ and $k\in\mathbb{Z}_{+}$.
As a corollary of the bound (1.5) above, in view of Lemma 2.19, there exists $s>0$ depending on $A,C$ and $q$ such that for any sequence $\{m_{n}\}_{n=1}^{\infty}$ satisfying $m_{n}\leq s\frac{n}{\log n}$, the total variation distance between $\mu_{n|m_{n}}$ and $\operatorname{Ber}_{m_{n}}\left(\frac{1}{1+C}\right)$ converges to zero as $n\rightarrow\infty$.
In the high density phase, as mentioned in the preface, the limiting measure is in general no longer product Bernoulli. Instead, the limiting measure will be characterized in the next definition in terms of the Askey–Wilson signed measures introduced in [30].
Definition 1.3.
In the high density phase $A>1$, $A>C$, we assume that $C/A\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $C\geq 1$. We define probability measures $\lambda_{m}$ on $\{0,1\}^{m}$ for $m\in\mathbb{Z}_{+}$ by their joint generating functions: For some $\varepsilon>0$ and for any $t_{1}\leq\dots\leq t_{m}$ within the interval $(1-\varepsilon,1]$,
$\mathbb{E}_{\lambda_{m}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]=\frac{A^{m}}{(1+A)^{2m}}\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\left(1+t_{i}+2\sqrt{t_{i}}x_{i}\right)\pi^{(A,1/A,C,D)}_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right),$ (1.6)
where $\pi_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)$ on the RHS above is the multi-dimensional Askey–Wilson signed measure (see Section 2.1 for a brief review). We will prove in Theorem 1.4 below that there exists $\varepsilon>0$ depending on $A,B,C,D$ and $q$ such that for all $m\in\mathbb{Z}_{+}$, the expression on the RHS above is indeed the generating function of a probability measure $\lambda_{m}$ on $\{0,1\}^{m}$, and that the marginal distribution of $\lambda_{m+1}$ on the leftmost sublattice $\{1,\dots,m\}$ coincides with $\lambda_{m}$. We define the probability measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ by requiring that its marginal distribution on the sublattice $\{1,\dots,m\}$ equals $\lambda_{m}$ for all $m=1,2,\dots$.
The next result is our main theorem in the high density phase.
Theorem 1.4.
In the high density phase $A>1$, $A>C$, we assume that $C/A\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $C\geq 1$. Then the probability measures $\lambda_{m}$ on $\{0,1\}^{m}$ for $m\in\mathbb{Z}_{+}$ and the probability measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ in Definition 1.3 are well-defined. Furthermore, as $n\rightarrow\infty$, the open ASEP stationary measure $\mu_{n}$ on $\{0,1\}^{n}$ weakly converges to the measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$.
Moreover, there exists $H>0$ depending on $A,B,C,D$ and $q$ and $\theta\in(0,1)$ depending on $A,C$ and $q$ such that for any $1\leq m\leq n$ we have:
$\mathsf{d}_{TV}\left(\mu_{n|m},\lambda_{m}\right)\leq\theta^{n}(Hm)^{3m},$ (1.7)
As a corollary of the bound (1.7) above, in view of Lemma 2.19, there exists $s>0$ depending on $A,C$ and $q$ such that for any sequence $\{m_{n}\}_{n=1}^{\infty}$ satisfying $m_{n}\leq s\frac{n}{\log n}$, the total variation distance between $\mu_{n|m_{n}}$ and $\lambda_{m_{n}}$ converges to zero as $n\rightarrow\infty$.
Theorem 1.2 and Theorem 1.4 above will be proved in Section 2.2.
Remark 1.5.
We note that Theorem 1.2 and Theorem 1.4, which respectively concern the low density and high density phases, are not related by the particle-hole duality (Lemma 2.7) of the open ASEP stationary measure. The particle-hole dual of Theorem 1.2 in the low density phase would correspond to a statement in the high density phase, but for the limit as $n\rightarrow\infty$ of a certain transformation of the open ASEP stationary measure—that is, the measure $\widetilde{\mu_{n}}$ on $\{0,1\}^{n}$ defined by $\widetilde{\mu_{n}}(\tau_{1},\dots,\tau_{n})=\mu_{n}(1-\tau_{n},\dots,1-\tau_{1})$ for any $\tau_{1},\dots,\tau_{n}\in\left\{0,1\right\}$, where $\mu_{n}$ is the open ASEP stationary measure. The two limiting measures on $\{0,1\}^{\mathbb{Z}_{+}}$ obtained respectively from $\widetilde{\mu_{n}}$ and $\mu_{n}$ on $\{0,1\}^{n}$ are fundamentally different.
Remark 1.6.
In our bounds (1.5) and (1.7) of the total variation distance, the constant $\theta\in(0,1)$ can be given by
$\theta=\frac{2+\max\left(2,qC+\left(qC\right)^{-1},\max\left(A,1\right)+\max\left(A,1\right)^{-1}\right)}{2+C+C^{-1}}$
in the low density phase and by the same formula with $A$ and $C$ swapped in the high density phase. The constant $s>0$ appearing in the growth rate $m_{n}\leq sn/\log n$ can be given by $s=-(\log\theta)/3$. These constants can be observed from our proofs of the main theorems in Section 2.2.
Remark 1.7.
Our total variation distance bounds (1.5) and (1.7) are not expected to be optimal. Hence the growth rate $n/\log n$ of sequences $\{m_{n}\}_{n=1}^{\infty}$ induced by those bounds are also not optimal. It would be an interesting question to ask for the optimal growth rate. Moreover, in the maximal current phase $A<1$, $C<1$, we do not know how to use our methods to obtain a total variation distance bound that is not too loose and that yields a growth rate of $\{m_{n}\}_{n=1}^{\infty}$ that is not too slow. Hence we do not cover this phase in the present paper. It remains an interesting question to find effective total variation distance bounds and the optimal growth rate of $\{m_{n}\}_{n=1}^{\infty}$ in the maximal current phase.
The next result determines exactly when the limiting measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ in the high density phase is a product Bernoulli measure.
Proposition 1.8.
Assume the same conditions as in Definition 1.3, i.e., we are in the high density phase $A>1$, $A>C$, and that $C/A\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $C\geq 1$. Then the measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ introduced in Definition 1.3 is a product Bernoulli measure if and only if $AC=1$, in which case it has density $A/(1+A)$.
Proposition 1.8 will be proved in Section 2.3. We note that when $AC=1$, it is known that the open ASEP stationary measure $\mu_{n}$ on $\{0,1\}^{n}$ is a product Bernoulli measure with density $A/(1+A)$, see for example [19, Appendix A]. Therefore the limiting measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ is also a product Bernoulli measure with the same density.
1.3 Comparision with a related result
As mentioned in the preface, a recent work by Nestoridi–Schmid [25] established a result which is of the same type as our Theorem 1.2 in the low density phase. We rephrase their result using our notation as follows.
Theorem 1.9 (Theorem 1.4 in [25]).
Assume that the open ASEP jump rates $\gamma=\delta=0$. Within the low density phase $C>A$, $C>1$ and the shock region $AC>1$, assume furthermore that there exist $\beta^{\prime}$ and $\beta^{\prime\prime}$ satisfying $\beta^{\prime}\leq\beta\leq\beta^{\prime\prime}$ and some $k\in\mathbb{Z}_{+}$ such that the respective parameters $A^{\prime}:=\phi_{+}(\beta^{\prime},\delta)$ and $A^{\prime\prime}:=\phi_{+}(\beta^{\prime\prime},\delta)$ (recall equations (1.2) and (1.3) defining $A,B,C,D$) satisfy $C>\max(A^{\prime},A^{\prime\prime},1)$ and $A^{\prime}Cq^{k}=A^{\prime\prime}Cq^{k-1}=1$. Then for any sequence $\{m_{n}\}_{n=1}^{\infty}\subset\mathbb{Z}_{+}$ satisfying $n-m_{n}\rightarrow\infty$ as $n\rightarrow\infty$,
$\lim_{n\rightarrow\infty}\mathsf{d}_{TV}\left(\mu_{n|m_{n}},\operatorname{Ber}_{m_{n}}\left(\frac{1}{1+C}\right)\right)=0.$
This theorem was proved in [25] using an entirely different method from the present paper. Their method relies on a characterization of the open ASEP stationary measures as convex linear combinations of specific Bernoulli shock measures, under the special condition $ACq^{k}=1$, as shown in Jafarpour–Masharian [22]. When comparing it with Theorem 1.2 in the low density phase, Theorem 1.9 (i.e., [25, Theorem 1.4]) provides a significantly faster growth rate for the sequence $\{m_{n}\}_{n=1}^{\infty}$, albeit within a smaller parameter range, assuming $\gamma=\delta=0$, and within a subregion of the shock region.
1.4 Outline of the paper
Section 2.1 reviews the Askey–Wilson signed measures and open ASEP stationary measures, along with their useful properties. In Section 2.2 we prove the main theorems using technical results that are provided in the three appendices. In Section 2.3 we prove Proposition 1.8. Appendix A establishes a bound of the total variation distance between two probability measures by their generating functions. Appendix B provides total variation bounds for certain Askey–Wilson signed measures. Appendix C proves a special symmetry of multi-dimensional Askey–Wilson signed measures known as the time reversal.
{acks}
[Acknowledgments] We thank Wlodek Bryc for bringing the paper [25] to our attention, for pointing out an error in the previous version of this manuscript, and for helpful discussions, which, in particular, contributed to Appendix C. We thank Ivan Corwin, Evita Nestoridi, Dominik Schmid, Yizao Wang and Jacek Wesołowski for helpful conversations. {funding} The author was partially supported by Ivan Corwin’s NSF grants DMS:1811143, DMS:2246576, Simons Foundation Grant 929852, and the Fernholz Foundation’s ‘Summer Minerva Fellows’ program.
2 Proofs of the main theorems
We review some background in Section 2.1 and prove the main theorems in Section 2.2. In Section 2.3 we will prove Proposition 1.8.
2.1 Background
We first review the definition and some results of Askey–Wilson signed measures following [30]. The Askey–Wilson signed measures were introduced in [30] generalizing the Askey–Wilson processes from the earlier works [8, 9]. We later review some useful properties of the open ASEP stationary measures.
Definition 2.1 (Definition 2.1 and Definition 2.2 in [30]).
Assume $q\in[0,1)$. We denote by $\Omega$ the set of parameters $(a,b,c,d)\in\mathbb{C}^{4}$ satisfying the following three assumptions:
  1. (1)
    $a,b$ are real, and $c,d$ are either real or form complex conjugate pair; $ab<1$ and $cd<1$,
  2. (2)
    for any two distinct $\mathfrak{e},\mathfrak{f}\in\{a,b,c,d\}$ such that $|\mathfrak{e}|,|\mathfrak{f}|\geq 1$, we have $\mathfrak{e}/\mathfrak{f}\notin\{q^{l}:l\in\mathbb{Z}\}$,
  3. (3)
    $q^{l}abcd\neq 1$ for all $l\in\mathbb{N}_{0}$, where $\mathbb{N}_{0}:=\{0,1,\dots\}$.
For $(a,b,c,d)\in\Omega$, the Askey–Wilson signed measure is of mixed type:
$\nu(\mathsf{d}x;a,b,c,d)=f(x;a,b,c,d)\mathds{1}_{|x|<1}\mathsf{d}x+\sum_{x\in F(a,b,c,d)}p(x)\delta_{x},$ (2.1)
where the continuous part density is defined as, for $x=\cos\theta\in(-1,1)$,
$f(x;a,b,c,d)=\frac{(q,ab,ac,ad,bc,bd,cd)_{\infty}}{2\pi(abcd)_{\infty}\sqrt{1-x^{2}}}\left\lvert\frac{(e^{2\mathtt{i}\theta})_{\infty}}{(ae^{\mathtt{i}\theta},be^{\mathtt{i}\theta},ce^{\mathtt{i}\theta},de^{\mathtt{i}\theta})_{\infty}}\right\lvert^{2}.$ (2.2)
Here and below, for complex $z$ and $n\in\mathbb{N}_{0}\cup\{\infty\}$, we use the $q$-Pochhammer symbol:
$(z)_{n}=(z;q)_{n}=\prod_{j=0}^{n-1}\,(1-zq^{j}),\quad(z_{1},\cdots,z_{k})_{n}=(z_{1},\cdots,z_{k};q)_{n}=\prod_{i=1}^{k}(z_{i};q)_{n}.$
The set of atoms $F(a,b,c,d)$ is generated by each $\mathfrak{e}\in\{a,b,c,d\}$ with $|\mathfrak{e}|\geq 1$. One can observe that the assumption $(a,b,c,d)\in\Omega$ guarantee that such $\mathfrak{e}$ is a real number. When $\mathfrak{e}=a$ the corresponding atoms are
$y_{k}^{\mathbf{a}}=y_{k}^{\mathbf{a}}(a,b,c,d)=\frac{1}{2}\left(aq^{k}+(aq^{k})^{-1}\right),$
with $k\geq 0$ such that $|aq^{k}|\geq 1$, and the corresponding masses are
$p(y_{0}^{\mathbf{a}})$ $=p_{0}^{\mathbf{a}}(a,b,c,d)=\frac{(a^{-2},bc,bd,cd)_{\infty}}{(b/a,c/a,d/a,abcd)_{\infty}},$ (2.3)
$p(y_{k}^{\mathbf{a}})$ $=p_{k}^{\mathbf{a}}(a,b,c,d)=\frac{p_{0}^{\mathbf{a}}(a,b,c,d)q^{k}(1-a^{2}q^{2k})(a^{2},ab,ac,ad)_{k}}{(q)_{k}(1-a^{2})a^{k}\prod_{l=1}^{k}\left((b-q^{l}a)(c-q^{l}a)(d-q^{l}a)\right)},\quad k\geq 1.$ (2.4)
The bold symbol $\mathbf{a}$ in the superscripts signal that the atom (if exists) is generated by the parameter coming from the $\mathbf{a}$ (first) position of the four parameters. For $\mathfrak{e}\in\{b,c,d\}$, atoms $y_{k}^{\mathbf{b}}$, $y_{k}^{\mathbf{c}}$, $y_{k}^{\mathbf{d}}$ and masses $p(y_{k}^{\mathbf{b}})$, $p(y_{k}^{\mathbf{c}})$, $p(y_{k}^{\mathbf{d}})$ are given by similar formulas with $a$ and $\mathfrak{e}$ swapped.
Remark 2.2.
Parameter $q\in[0,1)$ will be fixed throughout the paper.
Remark 2.3.
We recall from [30] that the Askey–Wilson signed measures are finite signed measures with compact supports in $\mathbb{R}$. Also they have total mass $1$, in the sense that:
$\int_{\mathbb{R}}\nu(\mathsf{d}x;a,b,c,d)=1.$
For certain parameters $A,B,C,D\in\mathbb{R}$ and on some suitable ‘time interval’ $I\subset\mathbb{R}$ (which will be defined later), we will study the following Askey–Wilson signed measures: For $t\in I$,
$\pi^{(A,B,C,D)}_{t}(\mathsf{d}y):=\nu\left(\mathsf{d}y;A\sqrt{t},B\sqrt{t},\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}}\right).$ (2.5)
Define $U_{t}\subset\mathbb{R}$ to be the (compact) support of $\pi^{(A,B,C,D)}_{t}(\mathsf{d}y)$. We will also study the following Askey–Wilson signed measures: For $s,t\in I$, $s<t$ and $x\in U_{s}$,
$P_{s,t}^{(A,B)}(x,\mathsf{d}y):=\nu\left(\mathsf{d}y;A\sqrt{t},B\sqrt{t},\sqrt{\frac{s}{t}}\left(x+\sqrt{x^{2}-1}\right),\sqrt{\frac{s}{t}}\left(x-\sqrt{x^{2}-1}\right)\right).$ (2.6)
When $s=t\in I$ and $x\in U_{s}$ we define $P^{(A,B)}_{s,s}(x,\mathsf{d}y)=\delta_{x}(\mathsf{d}y)$ for convenience.
For any $t_{1}\leq\dots\leq t_{n}$ in $I$, we define the ‘multi-dimensional’ Askey–Wilson signed measure:
$\pi^{(A,B,C,D)}_{t_{1},\dots,t_{n}}(\mathsf{d}x_{1},\dots,\mathsf{d}x_{n}):=\pi^{(A,B,C,D)}_{t_{1}}(\mathsf{d}x_{1})P^{(A,B)}_{t_{1},t_{2}}(x_{1},\mathsf{d}x_{2})\dots P^{(A,B)}_{t_{n-1},t_{n}}(x_{n-1},\mathsf{d}x_{n}),$ (2.7)
which is a finite signed measure with total mass $1$, supported on compact subset $U_{t_{1}}\times\dots\times U_{t_{n}}\subset\mathbb{R}^{n}$.
Remark 2.4.
We note that in this paper, the term “Askey–Wilson signed measure” is used to refer to several related objects: the signed measure $\nu(\mathsf{d}x;a,b,c,d)$ defined by (2.1); its parameter specializations $\pi_{t}^{(A,B,C,D)}(\mathsf{d}x)$ and $P_{s,t}^{(A,B)}(x,\mathsf{d}y)$ defined by (2.5) and (2.6) respectively; and the signed measure $\pi^{(A,B,C,D)}_{t_{1},\dots,t_{n}}(\mathsf{d}x_{1},\dots,\mathsf{d}x_{n})$ defined by (2.7), which we sometimes also refer to as the multi-dimensional Askey–Wilson signed measure.
In the rest of this paper, we will omit the superscripts $(A,B,C,D)$ and $(A,B)$ on the Askey–Wilson signed measures $\pi_{t}$, $\pi_{t_{1},\dots,t_{n}}$ and $P_{s,t}$ if no confusion will arise.
The following result characterizes the open ASEP stationary measures by Askey–Wilson signed measures. This characterization was originally due to [9] in the form of Askey–Wilson processes (see also anterior earlier work [29]) and was later generalized in [30].
Theorem 2.5.
Consider the open ASEP stationary measure $\mu_{n}=\mu_{n}^{(A,B,C,D)}$ on the lattice $\{1,\dots,n\}$. We assume that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exists $I=[1,1+\varepsilon)$ for some $\varepsilon=\varepsilon(A,B,C,q)>0$ depending on $A,B,C$ and $q$ such that for any $1\leq m<n$ and $t_{1}\leq\dots\leq t_{m}$ in $I$, we have:
$\mathbb{E}_{\mu_{n}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{n-m+i}}\right]\\ =\frac{\int_{\mathbb{R}^{m+1}}(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\pi_{1,t_{1},\dots,t_{m}}\left(\mathsf{d}x,\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)}{\int_{\mathbb{R}}(2+2x)^{n}\pi_{1}(\mathsf{d}x)}.$
We note that the formula above is the generating function for the open ASEP stationary measure $\mu_{n}$ on $\{0,1\}^{n}$, which involves the multi-dimensional Askey–Wilson signed measure parameterized by $(A,B,C,D)$. In contrast, formula (1.6) in Definition 1.3 gives the generating function for the limiting measure $\lambda_{m}$ on $\{0,1\}^{m}$, which involves the multi-dimensional Askey–Wilson signed measure parameterized by $(A,1/A,C,D)$.
Proof.
This theorem follows from [30, Theorem 1.1] except for the fact that $\varepsilon>0$ can be taken as depending on $A,B,C$ and $q$ but not on $D$. By the arguments in [30], we only need to show that there exists $\varepsilon=\varepsilon(A,B,C,q)>0$ such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$, we have $(a,b,c,d)\in\Omega$, where
$a=A\sqrt{t},\quad b=B\sqrt{t},\quad c=\sqrt{\frac{s}{t}}\left(x+\sqrt{x^{2}-1}\right),\quad d=\sqrt{\frac{s}{t}}\left(x-\sqrt{x^{2}-1}\right)$ (2.8)
so that $P_{s,t}^{(A,B)}(x,\mathsf{d}y)=\nu\left(\mathsf{d}y;a,b,c,d\right)$. Assumption (1) and (3) in Definition 2.1 always hold since $a$ and $b$ are real numbers; $c$ and $d$ form a complex conjugate pair; $ab=ABt\leq 0$; $cd=s/t<1$ and $abcd=ABs\leq 0$.
Recall that we have assumed $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. One can choose $\varepsilon=\varepsilon(A,B,C,q)>0$ such that $1+\varepsilon<\min\left(1/q,1/B^{2}\right)$ and $1+\varepsilon<1/A^{2}$ if $A<1$, and that the interval $I=[1,1+\varepsilon)$ does not contain elements in $\{Cq^{l}/A:l\in\mathbb{Z}\}$ if $A,C\geq 1$. In the following we will show that for any $s<t$ in $I$ and $x\in U_{s}$, assumption (2) in Definition 2.1 holds for $(a,b,c,d)$ given by (2.8). Note that we always have $b=B\sqrt{t}\in(-1,0]$ for $t\in I$. Also, since $B\sqrt{s},D/\sqrt{s}\in(-1,0]$, any possible atom in $U_{s}$ is generated by either $A\sqrt{s}$ or $C/\sqrt{s}$. We split the proof into three cases depending on $x\in U_{s}$:
  1. Case 1.
    Let $x\in[-1,1]$ then $c$ and $d$ are complex conjugate pairs with norm $<1$, where the norm of a complex number $z$ refers to its absolute value $|z|$. Hence at most one element in $\{a,b,c,d\}$ has norm $\geq 1$ and assumption (2) vacuously holds.
  2. Case 2.
    Let $x=\frac{1}{2}\left(q^{j}A\sqrt{s}+\left(q^{j}A\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}A\sqrt{s}>1$. Then $c=q^{j}As/\sqrt{t}$ and $d=1/\left(q^{j}A\sqrt{t}\right)<1$. Only $a$ and $c$ in $\{a,b,c,d\}$ can have norm $\geq 1$. We have $c/a=q^{j}s/t\notin\{q^{l}:l\in\mathbb{Z}\}$ since $s/t\in(q,1)$. Hence assumption (2) holds.
  3. Case 3.
    Let $x=\frac{1}{2}\left(q^{j}C/\sqrt{s}+\left(q^{j}C/\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}C/\sqrt{s}>1$. Then $c=q^{j}C/\sqrt{t}$ and $d=s/(q^{j}C\sqrt{t})<1$. Only $a$ and $c$ in $\{a,b,c,d\}$ can have norm $\geq 1$. If $A<1$ then $a=A\sqrt{t}<1$ and assumption (2) vacuously holds. If $C<1$ then $c=q^{j}C/\sqrt{t}<1$ and assumption (2) vacuously holds. If $A,C\geq 1$ then $c/a=q^{j}C/(At)\notin\{q^{l}:l\in\mathbb{Z}\}$ by our assumption, hence assumption (2) holds.
We conclude the proof. ∎
The next result is a basic property of the Askey–Wilson signed measure $P_{s,t}(x,\mathsf{d}y)$:
Lemma 2.6 (Lemma 2.14 and Lemma 2.15 in [30]).
Assume that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Choose the interval $I=[1,1+\varepsilon)$ from Theorem 2.5. Then for any $s\leq t$ in $I$ and any $x\in U_{s}$, the Askey–Wilson signed measure $P_{s,t}(x,\mathsf{d}y)$ is supported on $U_{t}$. In the high density phase, we have $P_{s,t}\left(y_{0}(s),\mathsf{d}y\right)=\delta_{y_{0}(t)}(\mathsf{d}y)$, where we denote by $y_{0}(t)$ the largest atom in $U_{t}$ for any $t\in I$.
We next recall some useful properties of the open ASEP stationary measure. The following result is well-known as the particle-hole duality:
Lemma 2.7 (Particle-hole duality).
The stationary measures $\mu_{n}^{(A,B,C,D)}$ and $\mu_{n}^{(C,D,A,B)}$ for open ASEP on $\{1,\dots,n\}$ are related by a transformation of occupation variables
[…]
See for example [30, Section 4.2] for the proof of the above result.
The open ASEP stationary measure is known to be continuous with respect to its parameters:
Lemma 2.8.
As a probability measure on $\{0,1\}^{n}$, the open ASEP stationary measure $\mu_{n}^{(A,B,C,D)}$ depends continuously on its parameters $A,B,C,D$ and $q$.
Proof.
It is shown in [2, Remark 1.9] that the open ASEP stationary measure depends real analytically on the jump rates $\alpha,\beta,\gamma,\delta$ and $q$ in the region of the parameter space where the stationary measure is unique. The continuity with respect to parameters $A,B,C,D$ then follows from [9, equation (2.4)]:
$\alpha=\frac{1-q}{(1+C)(1+D)},\quad\beta=\frac{1-q}{(1+A)(1+B)},\quad\\ \gamma=\frac{-(1-q)CD}{(1+C)(1+D)},\quad\delta=\frac{-(1-q)AB}{(1+A)(1+B)}.$
The following result is known as the ‘stochastic sandwiching’ of open ASEP stationary measure, which is first introduced by [12]:
Lemma 2.9 (Lemma 4.1 in [12]).
Fix $q\in[0,1)$ and consider real numbers
$0<\alpha^{\prime}\leq\alpha^{\prime\prime},\quad\beta^{\prime}\geq\beta^{\prime\prime}>0,\quad\gamma^{\prime}\geq\gamma^{\prime\prime}\geq 0,\quad 0\leq\delta^{\prime}\leq\delta^{\prime\prime}.$
Consider the open ASEP on the lattice $\{1,\dots,n\}$ with rates $(q,\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime})$ and $(q,\alpha^{\prime\prime},\beta^{\prime\prime},\gamma^{\prime\prime},\delta^{\prime\prime})$. We denote the stationary measures as $\mu_{n}^{\prime}$ and $\mu_{n}^{\prime\prime}$. The corresponding occupation variables are denoted as $(\tau^{\prime}_{1},\dots,\tau^{\prime}_{n})$ and $(\tau^{\prime\prime}_{1},\dots,\tau^{\prime\prime}_{n})$. Then there exists a coupling of $\mu^{\prime}_{n}$ and $\mu_{n}^{\prime\prime}$ such that almost surely $\tau^{\prime}_{i}\leq\tau_{i}^{\prime\prime}$ for $i=1,\dots,n$.
As a corollary, for any $t_{1},\dots,t_{n}\geq 1$, we have
$\mathbb{E}_{\mu^{\prime}_{n}}\left[\prod_{i=1}^{n}t_{i}^{\tau_{i}}\right]\leq\mathbb{E}_{\mu^{\prime\prime}_{n}}\left[\prod_{i=1}^{n}t_{i}^{\tau_{i}}\right].$
2.2 Proof of Theorem 1.2 and Theorem 1.4
In this subsection we will demonstrate the proofs of our main results: Theorem 1.2 in the low density phase and Theorem 1.4 in the high density phase.
For reasons that will become clear later, we first prove the results modulo a particle-hole duality. The main parts of these theorems (modulo the particle-hole duality) are stated in Theorem 2.11 below, which combines both the high and low density phases.
Definition 2.10.
Assume $\max(A,C)>1$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. We define probability measures $\eta_{m}$ on $\{0,1\}^{m}$ for $m\in\mathbb{Z}_{+}$ by their joint generating functions: For some $\varepsilon>0$ and for any $t_{1}\leq\dots\leq t_{m}$ within the interval $[1,1+\varepsilon)$,
$\mathbb{E}_{\eta_{m}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]\\ =\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\frac{1+t_{i}+2\sqrt{t_{i}}x_{i}}{2+2y_{0}(1)}P_{1,t_{1}}\left(y_{0}(1),\mathsf{d}x_{1}\right)P_{t_{1},t_{2}}\left(x_{1},\mathsf{d}x_{2}\right)\dots P_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right),$ (2.9)
where we recall from Lemma 2.6 that $y_{0}(1)$ is the largest atom in the (compact) support $U_{1}$ of the Askey–Wilson signed measure $\pi_{1}(\mathsf{d}y)=\nu(\mathsf{d}y;A,B,C,D)$, and $P_{s,t}(x,\mathsf{d}y)$ was defined by (2.6). We will prove in Theorem 2.11 that there exists $\varepsilon>0$ depending on $A,B,C,D$ and $q$ such that for all $m\in\mathbb{Z}_{+}$, the probability measure $\eta_{m}$ on $\{0,1\}^{m}$ are well-defined.
Theorem 2.11.
Assume $\max(A,C)>1$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then the probability measures $\eta_{m}$ on $\{0,1\}^{m}$ for $m\in\mathbb{Z}_{+}$ in Definition 2.10 are well-defined, and the marginal distribution of $\eta_{m+1}$ on the leftmost sublattice $\{1,\dots,m\}$ coincides with $\eta_{m}$.
Denote the marginal distribution of open ASEP stationary measure $\mu_{n}=\mu_{n}^{(A,B,C,D)}$ on the last $m$ sites $\{n-m+1,\dots,n\}$ by $\mu_{n,m}$. Then there exists $H>0$ depending on $A,B,C,D$ and $q$ and $\theta\in(0,1)$ depending on $A,C$ and $q$ such that for all $1\leq m\leq n$,
$\mathsf{d}_{TV}\left(\mu_{n,m},\eta_{m}\right)\leq\theta^{n}(Hm)^{3m}.$ (2.10)
Remark 2.12.
We note the difference between notations $\mu_{n|m}$ and $\mu_{n,m}$. The first notation denotes the marginal of $\mu_{n}$ on the first $m$ sites and the second one denotes the marginal on the last $m$ sites.
The proof of the above theorem will constitute the major technical component of this section. Before commencing with the proof, we will state two results which will be needed in the proof.
The following result provides a delicate bound of the total variation of Askey–Wilson signed measures $P_{s,t}(x,\mathsf{d}y)$:
Proposition 2.13.
Assume $A,C\geq 0$, $-1<B,D\leq 0$ and $q\in[0,1)$. Assume also that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exist positive constants $K\geq 1$ and $\varepsilon$ depending on $A,B,C$ and $q$ but not on $D$, such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$, the total variation of the Askey–Wilson signed measure $P_{s,t}(x,\mathsf{d}y)$ is bounded from above by $\frac{K}{(t-s)^{2}}$.
Remark 2.14.
The above result will appear again as Proposition B.1, which will be proved in Appendix B. It is worth noting that a strictly weaker version of this total variation bound has previously appeared in [30]. In particular, the total variation bound of $P_{s,t}(x,\mathrm{d}y)$ provided by [30, Proposition A.1] can be derived as a simple corollary of the above result, but the converse does not hold. The derivation of the bound here requires a delicate analysis of atom masses, see Appendix B.
To bound the total variation distance between two probability measures (in our case, $\mu_{n,m}$ and $\eta_{m}$) on $\{0,1\}^{m}$, we will use the following bound of this total variation distance by generating functions:
Proposition 2.15.
Let $\kappa$ and $\kappa^{\prime}$ be probability measures on $\{0,1\}^{m}$. Then for any set of numbers $0<t_{i,0}<t_{i,1}$ for $i=1,\dots,m$, we have:
$\mathsf{d}_{TV}(\kappa,\kappa^{\prime})\leq\frac{1}{2}\prod_{i=1}^{m}\frac{1+t_{i,1}}{t_{i,1}-t_{i,0}}\sum_{\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}}\left\lvert\mathbb{E}_{\kappa}\left[\prod_{i=1}^{m}t_{i,\upsilon_{i}}^{\tau_{i}}\right]-\mathbb{E}_{\kappa^{\prime}}\left[\prod_{i=1}^{m}t_{i,\upsilon_{i}}^{\tau_{i}}\right]\right\lvert,$
where $\tau_{i}\in\{0,1\}$ is the occupation variable on the site $i$, for $i=1,\dots,m$.
Remark 2.16.
The above result will appear again as Proposition A.1 which will be proved in Appendix A. As one can observe, this constitutes a general bound applicable to any two probability measures on $\{0,1\}^{m}$. While it is possible that this result is already known, we have been unable to locate it in references.
We now begin the proof of Theorem 2.11:
Proof of Theorem 2.11.
We divide the proof into three steps.
Step 1. In this step, under an extra assumption that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$, we prove that there exists $\varepsilon>0$ such that, for $1\leq t_{1}\leq\dots\leq t_{m}<1+\varepsilon$, we have:
$\lim_{n\rightarrow\infty}\mathbb{E}_{\mu_{m,n}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]\\ =\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\frac{1+t_{i}+2\sqrt{t_{i}}x_{i}}{2+2y_{0}(1)}P_{1,t_{1}}\left(y_{0}(1),\mathsf{d}x_{1}\right)P_{t_{1},t_{2}}\left(x_{1},\mathsf{d}x_{2}\right)\dots P_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right).$ (2.11)
The above equation in particular implies that its RHS (i.e., the RHS of (2.9)) indeed defines a probability measure $\eta_{m}$ on $\{0,1\}^{m}$ that is the limit of measures $\mu_{m,n}$ as $n\rightarrow\infty$. As a corollary, the marginal distribution of $\eta_{m+1}$ on the first $m$ sites coincides with $\eta_{m}$.
We choose $\varepsilon>0$ and $I=[1,1+\varepsilon)$ according to Theorem 2.5 and Proposition 2.13. By Theorem 2.5, we have that for any $1\leq m<n$ and $1=t_{0}\leq t_{1}\leq\dots\leq t_{m}<1+\varepsilon$,
$\mathbb{E}_{\mu_{n,m}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]=\mathbb{E}_{\mu_{n}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{n-m+i}}\right]=\frac{\Gamma_{1}^{\prime}}{\Gamma_{2}^{\prime}},$ (2.12)
where we define:
$\begin{split}\Gamma_{1}^{\prime}:=&\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\mathbb{R}^{m+1}}(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\pi_{1,t_{1},\dots,t_{m}}\left(\mathsf{d}x,\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right),\\ \Gamma_{2}^{\prime}:=&\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\mathbb{R}}(2+2x)^{n}\pi_{1}(\mathsf{d}x).\end{split}$ (2.13)
We write the RHS of (2.11) as $\Gamma_{1}/\Gamma_{2}$, where
$\begin{split}\Gamma_{1}:=&\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\frac{1+t_{i}+2\sqrt{t_{i}}x_{i}}{2+2y_{0}(1)}\pi_{1,t_{1},\dots,t_{m}}\left(\{y_{0}(1)\},\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)\\ =&\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\\ &\times\int_{\mathbb{R}^{m}}(2+2y_{0}(1))^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\pi_{1,t_{1},\dots,t_{m}}\left(\{y_{0}(1)\},\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right),\\ \Gamma_{2}:=&\pi_{1}\left(\{y_{0}(1)\}\right)=\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\{y_{0}(1)\}}(2+2x)^{n}\pi_{1}(\mathsf{d}x).\end{split}$ (2.14)
We need to bound the difference between $\Gamma_{1}/\Gamma_{2}$ and $\Gamma_{1}^{\prime}/\Gamma_{2}^{\prime}$. Under $|\Gamma_{2}|>|\Gamma_{2}^{\prime}-\Gamma_{2}|$, we have:
$\begin{split}\left\lvert\frac{\Gamma_{1}^{\prime}}{\Gamma_{2}^{\prime}}-\frac{\Gamma_{1}}{\Gamma_{2}}\right\rvert&=\frac{|\Gamma_{1}^{\prime}\Gamma_{2}-\Gamma_{1}\Gamma_{2}^{\prime}|}{|\Gamma_{2}^{\prime}\Gamma_{2}|}\leq\frac{|\Gamma_{1}^{\prime}\Gamma_{2}-\Gamma_{1}\Gamma_{2}|+|\Gamma_{1}\Gamma_{2}-\Gamma_{1}\Gamma_{2}^{\prime}|}{|\Gamma_{2}^{\prime}\Gamma_{2}|}\\ &=\frac{|\Gamma_{1}^{\prime}-\Gamma_{1}|}{|\Gamma_{2}^{\prime}|}+\frac{|\Gamma_{1}|}{|\Gamma_{2}^{\prime}\Gamma_{2}|}|\Gamma_{2}^{\prime}-\Gamma_{2}|\leq\frac{|\Gamma_{1}^{\prime}-\Gamma_{1}|}{|\Gamma_{2}|-|\Gamma_{2}^{\prime}-\Gamma_{2}|}+\frac{|\Gamma_{1}|}{|\Gamma_{2}|}\frac{|\Gamma_{2}^{\prime}-\Gamma_{2}|}{|\Gamma_{2}|-|\Gamma_{2}^{\prime}-\Gamma_{2}|}.\end{split}$ (2.15)
Hence we need to bound three quantities $|\Gamma_{1}/\Gamma_{2}|$, $|\Gamma_{1}^{\prime}-\Gamma_{1}|$ and $|\Gamma_{2}^{\prime}-\Gamma_{2}|$. We first recall that $\Gamma_{1}/\Gamma_{2}$ is the RHS of (2.11), i.e.,
$\frac{\Gamma_{1}}{\Gamma_{2}}=\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\frac{1+t_{i}+2\sqrt{t_{i}}x_{i}}{2+2y_{0}(1)}P_{1,t_{1}}\left(y_{0}(1),\mathsf{d}x_{1}\right)P_{t_{1},t_{2}}\left(x_{1},\mathsf{d}x_{2}\right)\dots P_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right).$ (2.16)
In view of Lemma 2.6, the signed measure
$P_{1,t_{1}}\left(y_{0}(1),\mathsf{d}x_{1}\right)\dots P_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right)$
on the RHS is supported on $U_{t_{1}}\times\dots\times U_{t_{m}}$. We next estimate its total variation. We denote the total variation of any signed measure $\nu$ by $\left\lVert\nu\right\rVert_{TV}$. Notice that for $1\leq i\leq m$ and $x\in U_{t_{i-1}}$, we have $\left\lVert P_{t_{i-1},t_{i}}\left(x,\mathsf{d}y\right)\right\rVert_{TV}=1$ when $t_{i-1}=t_{i}$ and, by Proposition 2.13, $\left\lVert P_{t_{i-1},t_{i}}\left(x,\mathsf{d}y\right)\right\rVert_{TV}\leq\frac{K}{(t_{i}-t_{i-1})^{2}}$ when $t_{i-1}<t_{i}$. Therefore, by multiplying those bounds together, we have:
$\left\lVert P_{1,t_{1}}\left(y_{0}(1),\mathsf{d}x_{1}\right)P_{t_{1},t_{2}}\left(x_{1},\mathsf{d}x_{2}\right)\dots P_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right)\right\rVert_{TV}\leq\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}}.$ (2.17)
Notice that $01+t+2\sqrt{t}x>$ 0 for any $x\in U_{t}$. Define
$r:=\sup_{t\in I}\left(1+t+2\sqrt{t}y_{0}(t)\right)$
In view of (2.16) and (2.17), and noting that $12+2y_{0}(t)\geq 4>$ 1, we have:
$\left\lvert\frac{\Gamma_{1}}{\Gamma_{2}}\right\lvert\leq\left(\frac{r}{2+2y_{0}(1)}\right)^{m}\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}}\leq r^{m}\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}}.$ (2.18)
We next bound the differences $|\Gamma_{1}^{\prime}-\Gamma_{1}|$ and $|\Gamma_{2}^{\prime}-\Gamma_{2}|$. By (2.13) and (2.14) we have:
$\Gamma_{1}^{\prime}-\Gamma_{1}=\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{\left(U_{1}\setminus\{y_{0}(1)\}\right)\times U_{t_{1}}\times\dots\times U_{t_{m}}}(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})\\ \times\pi_{1,t_{1},\dots,t_{m}}\left(\mathsf{d}x,\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right).$ (2.19)
Using Proposition 2.13 in a similar way as above, we have:
$\left\lVert\pi_{1,t_{1},\dots,t_{m}}\right\rVert_{TV}\leq\left\lVert\pi_{1}\right\rVert_{TV}\prod_{i=1}^{m}\sup_{x\in U_{t_{i-1}}}\left\lVert P_{t_{i-1},t_{i}}\left(x,\mathsf{d}y\right)\right\rVert_{TV}\\ \leq\left\lVert\pi_{1}\right\rVert_{TV}\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}}.$
For $t\in I$, we denote by $y_{0}^{*}(t)$ the largest element of $U_{t}\setminus\{y_{0}(t)\}$. In the support
$\left(U_{1}\setminus\{y_{0}(1)\}\right)\times U_{t_{1}}\times\dots\times U_{t_{m}}$
on the RHS integration of (2.19), we have $x\leq y^{*}_{0}(1)$ and $x_{i}\leq y_{0}(t_{i})$ for $1\leq i\leq m$. Hence:
$0<\frac{(2+2x)^{n-m}\prod_{i=1}^{m}(1+t_{i}+2\sqrt{t_{i}}x_{i})}{\left(2+2y_{0}(1)\right)^{n}}\\ \leq\frac{(2+2y^{*}_{0}(1))^{n-m}r^{m}}{\left(2+2y_{0}(1)\right)^{n}}<\frac{(2+2y^{*}_{0}(1))^{n}r^{m}}{\left(2+2y_{0}(1)\right)^{n}}=\theta^{n}r^{m},$
where we denote $\theta:=\frac{1+y^{*}_{0}(1)}{1+y_{0}(1)}\in(0,1)$. Therefore, in view of (2.19), we have
$|\Gamma_{1}^{\prime}-\Gamma_{1}|\leq\theta^{n}r^{m}\left\lVert\pi_{1}\right\rVert_{TV}\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}}.$ (2.20)
By (2.13) and (2.14) we also have:
$\Gamma_{2}^{\prime}-\Gamma_{2}=\frac{1}{\left(2+2y_{0}(1)\right)^{n}}\int_{U_{1}\setminus\{y_{0}(1)\}}(2+2x)^{n}\pi_{1}(\mathsf{d}x).$
Using a similar but simpler argument to the one above, we conclude
$|\Gamma_{2}^{\prime}-\Gamma_{2}|\leq\theta^{n}\left\lVert\pi_{1}\right\rVert_{TV}.$
We recall that $\Gamma_{2}=\pi_{1}\left(\{y_{0}(1)\}\right)$. There exists $N=N(A,B,C,D,q)\in\mathbb{Z}_{+}$ such that for any $n\geq N$, we have
$|\Gamma_{2}|=\left\lvert\pi_{1}\left(\{y_{0}(1)\}\right)\right\lvert>\theta^{n}\left\lVert\pi_{1}\right\rVert_{TV}\geq|\Gamma_{2}^{\prime}-\Gamma_{2}|.$ (2.21)
Combining the estimates (2.15), (2.18), (2.20) and (2.21), we get that for any $n\geq N$, $1\leq m<n$ and $1=t_{0}\leq t_{1}\leq\dots\leq t_{m}<1+\varepsilon$,
$\left\lvert\frac{\Gamma^{\prime}_{1}}{\Gamma^{\prime}_{2}}-\frac{\Gamma_{1}}{\Gamma_{2}}\right\rvert\leq\frac{2\theta^{n}r^{m}\left\lVert\pi_{1}\right\rVert_{TV}}{\left\lvert\pi_{1}\left(\{y_{0}(1)\}\right)\right\lvert-\theta^{n}\left\lVert\pi_{1}\right\rVert_{TV}}\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}}.$ (2.22)
For fixed $m$, as $n\rightarrow\infty$, the RHS above converges to 00. In view of (2.12) and the definitions of $\Gamma_{1}$ and $\Gamma_{2}$ (so that the RHS of (2.11) equals $\Gamma_{1}/\Gamma_{2}$), we conclude the proof of the convergence (2.11). Step 1 is complete.
Step 2. In this step, under the extra assumption that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$, we prove (2.10): there exists $H>0$ depending on $A,B,C,D$ and $q$ such that for all $1\leq m\leq n$,
$\mathsf{d}_{TV}\left(\mu_{n,m},\eta_{m}\right)\leq\theta^{n}(Hm)^{3m},$ (2.23)
where we recall $\theta=\frac{1+y^{*}_{0}(1)}{1+y_{0}(1)}\in(0,1)$.
To bound the total variation distance of two probability measures on $\{0,1\}^{m}$, by Proposition 2.15, we only need to bound the difference between their generating functions. In view of (2.22), we have:
$\left\lvert\mathbb{E}_{\mu_{n,m}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]-\mathbb{E}_{\eta_{m}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]\right\rvert\leq\frac{2\theta^{n}r^{m}\left\lVert\pi_{1}\right\rVert_{TV}}{\left\lvert\pi_{1}\left(\{y_{0}(1)\}\right)\right\lvert-\theta^{n}\left\lVert\pi_{1}\right\rVert_{TV}}\prod_{1\leq i\leq m,t_{i-1}<t_{i}}\frac{K}{(t_{i}-t_{i-1})^{2}},$ (2.24)
for any $n\geq N$, $1\leq m<n$ and $1=t_{0}\leq t_{1}\leq\dots\leq t_{m}<1+\varepsilon$. We set
$t_{i,\upsilon,m}:=1+\frac{(i+\upsilon)\varepsilon}{2m}\quad\text{for }i=1,\dots,m\text{ and }\upsilon\in\{0,1\}.$
For any $\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}$, we take $t_{i}=t_{i,\upsilon_{i},m}$ for $i=1,\dots,m$ in (2.24). In view of the fact that
$t_{i}-t_{i-1}\in\left\{0,\frac{\varepsilon}{2m},\frac{\varepsilon}{m}\right\},\quad i=1,\dots,m,$
we have:
$\left\lvert\mathbb{E}_{\mu_{n,m}}\left[\prod_{i=1}^{m}t_{i,\upsilon_{i},m}^{\tau_{i}}\right]-\mathbb{E}_{\eta_{m}}\left[\prod_{i=1}^{m}t_{i,\upsilon_{i},m}^{\tau_{i}}\right]\right\rvert\leq\frac{2\theta^{n}r^{m}\left\lVert\pi_{1}\right\rVert_{TV}}{\left\lvert\pi_{1}\left(\{y_{0}(1)\}\right)\right\lvert-\theta^{n}\left\lVert\pi_{1}\right\rVert_{TV}}\frac{K^{m}}{(\varepsilon/(2m))^{2m}}.$ (2.25)
Using Proposition 2.15, in view of
$1+t_{i,1,m}<2+\varepsilon,\quad t_{i,1,m}-t_{i,0,m}=\frac{\varepsilon}{2m},\quad i=1,\dots,m,$
we have:
$\mathsf{d}_{TV}\left(\mu_{n,m},\eta_{m}\right)\leq\frac{\theta^{n}r^{m}\left\lVert\pi_{1}\right\rVert_{TV}}{\left\lvert\pi_{1}\left(\{y_{0}(1)\}\right)\right\lvert-\theta^{n}\left\lVert\pi_{1}\right\rVert_{TV}}\frac{K^{m}}{(\varepsilon/(2m))^{2m}}\left(\frac{2+\varepsilon}{\varepsilon/(2m)}\right)^{m}2^{m}.$ (2.26)
One can choose $H=H(A,B,C,D,q)>0$ sufficiently large so that the RHS of (2.26) can be bounded above by $\theta^{n}(Hm)^{3m}$ for any $n\geq N$ and $1\leq m<n$. For the cases when $n<N$ or $m=n$, one can increase the value of $H$ if necessary, so that $\theta^{n}(Hm)^{3m}\geq 1$, which is greater or equal to $\mathsf{d}_{TV}\left(\mu_{n,m},\eta_{m}\right)\leq 1$. We conclude the proof of (2.23). Step 2 is complete.
Step 3. In this step we show that the total variation distance bound (2.23) continues to hold without the assumption that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$. Note that the well-definedness of the measure $\eta_{m}$ (which we concluded from Step 1) does not have any issues, since it only involves $A,B,C$ and $q$.
Before we embark on the major component of Step 3, we present the following technical total variation bound of the Askey–Wilson signed measure $\pi_{1}$ by the mass of its largest atom:
Proposition 2.17.
Assume $A,C\geq 0$, $\max(A,C)>1$, $-1<B,D\leq 0$, $q\in[0,1)$ and $|(ABCD)_{\infty}|\leq 1$. Assume also $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exists a positive constant $L$ depending on $A,C$ and $q$ such that
$\left\lVert\pi_{1}\right\rVert_{TV}\leq\frac{L}{1-BD}\left\lvert\pi_{1}\left(\{y_{0}(1)\}\right)\right\lvert.$
Remark 2.18.
The above result will appear again as Proposition B.6 which will be proved in Appendix B by bounding the norms of atom masses in this Askey–Wilson signed measure.
We return to Step 3 in the proof. We consider the case $q^{j}ABCD=1$ for some $j\in\mathbb{N}_{0}$. We will use the continuity argument. We take a sequence of $D_{k}\in(-1,0)$, $k=1,2,\dots$ that converges to $D$, satisfying $q^{l}ABCD_{k}\neq 1$ for all $l\in\mathbb{N}_{0}$ and for all $k$. Since $(ABCD)_{\infty}=0$, we can assume that $|(ABCD_{k})_{\infty}|\leq 1$ for all $k$. Using the continuity of the stationary measure $\mu_{n}^{(A,B,C,D)}$ from Lemma 2.8, we have:
$\lim_{k\rightarrow\infty}\mathsf{d}_{TV}\left(\mu^{(A,B,C,D_{k})}_{n,m},\eta_{m}\right)=\mathsf{d}_{TV}\left(\mu^{(A,B,C,D)}_{n,m},\eta_{m}\right).$ (2.27)
Using the conclusion of Step 2, for each $k=1,2,\dots$ one can bound $\mathsf{d}_{TV}\left(\mu^{(A,B,C,D_{k})}_{n,m},\eta_{m}\right)$ by the corresponding RHS of (2.26). We observe that most of the terms therein do not rely on $k$:
  1. $\bullet$
    The values of $\varepsilon$ and $K$ were selected by Proposition 2.13, which only involve $A,B,C$ and $q$ and hence do not rely on $k$.
  2. $\bullet$
    The values of $\theta=\frac{1+y^{*}_{0}(1)}{1+y_{0}(1)}$ and $r=\sup_{t\in I}\left(1+t+2\sqrt{t}y_{0}(t)\right)$ depend solely on $\varepsilon,A,C$ and $q$ (recall that $I=[1,1+\varepsilon)$) and hence do not rely on $k$.
Therefore, in view of (2.27), we have:
$\begin{split}\mathsf{d}_{TV}\left(\mu^{(A,B,C,D)}_{n,m},\eta_{m}\right)\leq&\theta^{n}r^{m}\frac{K^{m}}{(\varepsilon/(2m))^{2m}}\left(\frac{2+\varepsilon}{\varepsilon/(2m)}\right)^{m}2^{m}\\ &\times\limsup_{k\rightarrow\infty}\frac{\left\lVert\pi^{(A,B,C,D_{k})}_{1}\right\rVert_{TV}}{\left\lvert\pi^{(A,B,C,D_{k})}_{1}\left(\{y_{0}(1)\}\right)\right\lvert-\theta^{n}\left\lVert\pi^{(A,B,C,D_{k})}_{1}\right\rVert_{TV}}.\end{split}$
Using Proposition 2.17 above, in view of our assumption $|(ABCD_{k})_{\infty}|\leq 1$ and the fact that $1/(1-BD_{k})$ is bounded by a uniform constant for all $k=1,2,\dots$, we conclude that there exists $N$ independent of $k$ such that for all $n\geq N$, the supremum limit in the RHS above is a finite positive number. Therefore the bound (2.23):
$\mathsf{d}_{TV}\left(\mu_{n,m},\eta_{m}\right)\leq\theta^{n}(Hm)^{3m}$
holds at $(A,B,C,D)$ for some positive number $H=H(A,B,C,D)$ for $n\geq N$ and $1\leq m<n$. Similar to the end of Step 2, for $n<N$ or $m=n$, one can increase the value of $H$ if necessary so that $\mathsf{d}_{TV}\left(\mu_{n,m},\eta_{m}\right)\leq 1$ can still be bounded by $\theta^{n}(Hm)^{3m}$. Step 3 is complete.
In view of Step 1, Step 2 and Step 3 above, we conclude the proof of Theorem 2.11. ∎
In the rest of this subsection we will use Theorem 2.11 and the particle-hole duality to deduce Theorem 1.2 and Theorem 1.4. We recall the marginal distributions $\mu_{n|m}$ and $\mu_{n,m}$ of $\mu_{n}$ and Remark 2.12. In view of the particle-hole duality stated as Lemma 2.7, we have:
$\mu_{n|m}^{(A,B,C,D)}(\tau_{1},\dots,\tau_{m})=\mu_{n,m}^{(C,D,A,B)}(\tau_{m},\dots,\tau_{1})\quad\text{for any }\tau_{1},\dots,\tau_{m}\in\{0,1\}.$ (2.28)
The next technical lemma shows that the total variation distance bound $\theta^{n}(Hm)^{3m}$ induces the correct growth rate of $\{m_{n}\}_{n=1}^{\infty}$. Note that we only need $p=3$ in this lemma.
Lemma 2.19.
For any $\theta\in(0,1)$, $H>0$ and $p>0$, we denote $s=-\frac{1}{p}\log\theta>0$. Then for any sequence $\{m_{n}\}_{n=1}^{\infty}$ satisfying $1\leq m_{n}\leq s\frac{n}{\log n}$ for $n=1,2,\dots$, we have
$\lim_{n\rightarrow\infty}\theta^{n}(Hm_{n})^{pm_{n}}=0.$
Proof.
For any constant $R>0$, for sufficiently large $n$, we have:
$(Rm_{n})^{pm_{n}}\leq\left(\frac{Rsn}{\log n}\right)^{\frac{psn}{\log n}}=\left(\frac{Rs}{\log n}n\right)^{\frac{psn}{\log n}}\leq n^{\frac{psn}{\log n}}\\ =n^{\frac{n\log\left(\frac{1}{\theta}\right)}{\log n}}=\left(n^{\frac{\log\left(\frac{1}{\theta}\right)}{\log n}}\right)^{n}=\left(n^{\log_{n}\left(\frac{1}{\theta}\right)}\right)^{n}=\left(\frac{1}{\theta}\right)^{n},$
hence
$\overline{\lim}_{n\rightarrow\infty}\theta^{n}(Hm_{n})^{pm_{n}}\leq\overline{\lim}_{n\rightarrow\infty}(H/R)^{pm_{n}}\leq(H/R)^{p}.$
Taking $R\rightarrow\infty$ we conclude the proof. ∎
We now prove the main Theorem 1.2 in the low density phase.
Proof of Theorem 1.2.
Although this result is about the low density phase, we will use Theorem 2.11 in the high density phase and later use particle-hole duality.
In the high density phase $A>1$, $A>C$, we assume $C/A\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $C\geq 1$ and analyze the measure $\eta_{m}$ introduced in Definition 2.10. In view of the second statement in Lemma 2.6, we have
$P_{1,t_{1}}\left(y_{0}(1),\mathsf{d}x_{1}\right)\dots P_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right)=\delta_{y_{0}(t_{1})}(\mathsf{d}x_{1})\dots\delta_{y_{0}(t_{m})}(\mathsf{d}x_{m}).$
We notice $y_{0}(t)=\frac{1}{2}\left(A\sqrt{t}+\frac{1}{A\sqrt{t}}\right)$ and therefore
$\frac{1+t_{i}+2\sqrt{t_{i}}y_{0}(t)}{2+2y_{0}(1)}=\frac{1+At_{i}}{1+A}.$
Hence the RHS of (2.9) equals $\prod_{i=1}^{n}\frac{1+At_{i}}{1+A}$, using which we conclude that $\eta_{m}=\operatorname{Ber}_{m}\left(\frac{A}{1+A}\right)$.
For the total variation distance bound, by Theorem 2.11 we have
$\mathsf{d}_{TV}\left(\mu_{n,m},\operatorname{Ber}_{m}\left(\frac{A}{1+A}\right)\right)\leq\theta^{n}(Hm)^{3m}.$
In view of particle-hole duality (2.28), in the low density phase $C>1$, $C>A$, assuming that $A/C\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $A\geq 1$, we have that for any $1\leq m\leq n$,
$\mathsf{d}_{TV}\left(\mu_{n|m},\operatorname{Ber}_{m}\left(\frac{1}{1+C}\right)\right)\leq\theta^{n}(Hm)^{3m}.$ (2.29)
We conclude the total variation bound (1.5) in Theorem 1.2. The last statement in the theorem (about the sequence $\{m_{n}\}$) follows directly from this bound combining with technical Lemma 2.19.
For fixed $m$, we note that the RHS of (2.29) converges to 00 as $n\rightarrow\infty$. In particular, the sequence of measures $\mu_{n|m}$ converges to $\operatorname{Ber}_{m}\left(\frac{1}{1+C}\right)$. Next we show this convergence holds true without the assumption $A/C\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ for $A\geq 1$. We fix the jump rates $\alpha,\beta,\gamma,q$ and choose $\delta^{\prime}>\delta$ so that $A^{\prime}:=\phi_{+}(\beta,\delta^{\prime})$ satisfy $A^{\prime}/C\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$. In view of the stochastic sandwiching Lemma 2.9, we have
$\mathbb{E}_{\mu_{n|m}^{(A^{\prime},B,C,D)}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]\geq\mathbb{E}_{\mu_{n|m}^{(A,B,C,D)}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right],$ (2.30)
for any $t_{1},\dots,t_{m}\geq 1$. As $n\rightarrow\infty$, we know that $\mu_{n|m}^{(A^{\prime},B,C,D)}$ converges to $\operatorname{Ber}_{m}\left(\frac{1}{1+C}\right)$, therefore the LHS of (2.30) converges to
$\mathbb{E}_{\operatorname{Ber}_{m}\left(\frac{1}{1+C}\right)}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]=\prod_{i=1}^{m}\frac{t_{i}+C}{1+C}.$
We conclude
$\prod_{i=1}^{m}\frac{t_{i}+C}{1+C}\geq\limsup_{n\rightarrow\infty}\mathbb{E}_{\mu_{n|m}^{(A,B,C,D)}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right].$
By the same reason, the opposite inequality holds true. Therefore the convergence holds true for $(A,B,C,D)$. We conclude that on the entire high density phase, the sequence of measures $\mu_{n}$ converges weakly to the product Bernoulli measure with density $\frac{1}{1+C}$ on $\{0,1\}^{\mathbb{Z}_{+}}$. The proof is concluded. ∎
Before we provide the proof of Theorem 1.4, we present a special symmetry of multi-dimensional Askey–Wilson signed measures known as the ‘time reversal’, which will be needed in the proof.
Proposition 2.20.
Assume $q\in[0,1)$ and $A,B,C,D\in\mathbb{R}$. We have:
$\pi_{t_{1},\dots,t_{m}}^{(A,B,C,D)}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)=\pi_{1/t_{m},\dots,1/t_{1}}^{(C,D,A,B)}\left(\mathsf{d}x_{m},\dots,\mathsf{d}x_{1}\right)$ (2.31)
for any $0<t_{1}\leq\dots\leq t_{m}$ for which the multi-dimensional Askey–Wilson signed measures on both sides of this identity are well-defined.
Remark 2.21.
The above result will reappear as Proposition C.1 and will be proved in Appendix C.
We now prove the main Theorem 1.4 in the high density phase.
Proof of Theorem 1.4.
Although this result is about the high density phase, we will use Theorem 2.11 in the low density phase and later use particle-hole duality to conclude the proof.
In the low density phase $C>1$, $C>A$, we assume $A/C\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $A\geq 1$ and analyze the measure $\eta_{m}$ introduced in Definition 2.10. Notice $y_{0}(t)=\frac{1}{2}\left(\frac{C}{\sqrt{t}}+\frac{\sqrt{t}}{C}\right)$ and hence
$P_{1,t_{1}}^{(A,B)}\left(y_{0}(1),\mathsf{d}x_{1}\right)=\nu\left(\mathsf{d}x_{1};A\sqrt{t_{1}},B\sqrt{t_{1}},C/\sqrt{t_{1}},1/\left(C\sqrt{t_{1}}\right)\right)=\pi_{t_{1}}^{(A,B,C,1/C)}(\mathsf{d}x_{1}).$
Therefore the signed measure on the RHS of (2.9) equals:
$P_{1,t_{1}}^{(A,B)}\left(y_{0}(1),\mathsf{d}x_{1}\right)\dots P^{(A,B)}_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right)\\ =\pi_{t_{1}}^{(A,B,C,1/C)}(\mathsf{d}x_{1})P^{(A,B)}_{t_{1},t_{2}}\left(x_{1},\mathsf{d}x_{2}\right)\dots P^{(A,B)}_{t_{m-1},t_{m}}\left(x_{m-1},\mathsf{d}x_{m}\right)\\ =\pi^{(A,B,C,1/C)}_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right),$
Since $2+2y_{0}(1)=(1+C)^{2}/C$, by (2.9) we conclude that, for any $1\leq t_{1}\leq\dots\leq t_{m}<1+\varepsilon$,
$\mathbb{E}_{\eta_{m}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]=\frac{C^{m}}{(1+C)^{2m}}\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\left(1+t_{i}+2\sqrt{t_{i}}x_{i}\right)\pi^{(A,B,C,1/C)}_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right).$ (2.32)
We define the measure $\lambda_{m}$ (whose existence is claimed in Theorem 1.4) as the particle-hole duality of $\eta_{m}$, i.e. in the high density phase $A>1$, $A>C$, assuming $C/A\notin\{q^{l}:l\in\mathbb{Z}_{+}\}$ if $C\geq 1$, we define:
$\lambda_{m}^{(A,B,C,D)}(\tau_{1},\dots,\tau_{m})=\eta_{m}^{(C,D,A,B)}(1-\tau_{m},\dots,1-\tau_{1})\quad\text{for all}\quad\tau_{1},\dots,\tau_{m}\in\{0,1\}.$
Therefore, for any $\frac{1}{1+\varepsilon}<t_{1}\leq\dots\leq t_{m}\leq 1$, we have:
$\begin{split}&\mathbb{E}_{\lambda_{m}^{(A,B,C,D)}}\left[\prod_{i=1}^{m}t_{i}^{\tau_{i}}\right]=\mathbb{E}_{\eta_{m}^{(C,D,A,B)}}\left[\prod_{i=1}^{m}t_{m+1-i}^{1-\tau_{i}}\right]\\ &=t_{1}\dots t_{m}\mathbb{E}_{\eta_{m}^{(C,D,A,B)}}\left[\prod_{i=1}^{m}\left(\frac{1}{t_{m+1-i}}\right)^{\tau_{i}}\right]\\ &=t_{1}\dots t_{m}\frac{A^{m}}{(1+A)^{2m}}\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\left(1+\frac{1}{t_{m+1-i}}+2\sqrt{\frac{1}{t_{m+1-i}}}x_{i}\right)\pi^{(C,D,A,1/A)}_{\frac{1}{t_{m}},\dots,\frac{1}{t_{1}}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)\\ &=\frac{A^{m}}{(1+A)^{2m}}\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\left(1+t_{m+1-i}+2\sqrt{t_{m+1-i}}x_{i}\right)\pi^{(A,1/A,C,D)}_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{m},\dots,\mathsf{d}x_{1}\right)\\ &=\frac{A^{m}}{(1+A)^{2m}}\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\left(1+t_{i}+2\sqrt{t_{i}}x_{i}\right)\pi^{(A,1/A,C,D)}_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right),\end{split}$
where we have used (2.32) for $1\leq 1/t_{m}\leq\dots\leq 1/t_{1}<1+\varepsilon$ and the time reversal symmetry of Askey–Wilson signed measures (Proposition 2.20). The above identity coincides with (1.6) in Definition 1.3. By particle-hole duality, the total variation distance bound (1.7) in Theorem 1.4 directly follows from the bound (2.10) in Theorem 2.11. The last statement in the theorem (about the sequence $\{m_{n}\}$) follows directly from this bound combining with the technical Lemma 2.19. We conclude the proof. ∎
2.3 Proof of Proposition 1.8
In this subsection we prove Proposition 1.8, which states that the limiting measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ in the high density phase, as introduced in Definition 1.3, is a product Bernoulli measure if and only if $AC=1$, in which case it has density $A/(1+A)$.
Proof of Proposition 1.8.
Assume that the measure $\lambda$ on $\{0,1\}^{\mathbb{Z}_{+}}$ is product Bernoulli with density $\rho\in[0,1]$, then each $\lambda_{m}$ on $\{0,1\}^{m}$ for $m\in\mathbb{Z}_{+}$ is product Bernoulli with density $\rho$. Take $t_{1}=\dots=t_{m}=t\in(1-\varepsilon,1)$ in (1.6) we have
$\begin{split}\mathbb{E}_{\lambda_{m}}\left[t^{\tau_{1}+\dots+\tau_{m}}\right]&=(t\rho+1-\rho)^{m}\\ &=\frac{A^{m}}{(1+A)^{2m}}\int_{\mathbb{R}^{m}}\prod_{i=1}^{m}\left(1+t+2\sqrt{t}x_{i}\right)\pi^{(A,1/A,C,D)}_{t,\dots,t}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)\\ &=\frac{A^{m}}{(1+A)^{2m}}\int_{\mathbb{R}}\left(1+t+2\sqrt{t}x\right)^{m}\pi_{t}^{(A,1/A,C,D)}(\mathsf{d}x).\end{split}$ (2.33)
Recall that $\pi^{(A,1/A,C,D)}_{t}(\mathsf{d}x)=\nu\left(\mathsf{d}x;A\sqrt{t},\sqrt{t}/A,C/\sqrt{t},D/\sqrt{t}\right)$. Since $A>C\geq 0$, $A>1$ and $D\in(-1,0]$, one can shrink the value of $\varepsilon>0$ if necessary, so that for all $t\in(1-\varepsilon,1)$, we have $A\sqrt{t}>1$, $A\sqrt{t}>C/\sqrt{t}\geq 0$, $\sqrt{t}/A\in[0,1)$ and $D/\sqrt{t}\in(-1,0]$. Therefore the signed measure $\pi_{t}(\mathsf{d}x)$ only has atoms $>1$, and the largest atom is $y_{0}(t)=\frac{1}{2}\left(A\sqrt{t}+\frac{1}{A\sqrt{t}}\right)$. Using formula (2.3) one can check that $\pi_{t}(\{y_{0}(t)\})\neq 0$. We denote by $y_{1}^{*}(t)$ the second largest atom of $\pi_{t}(\mathsf{d}x)$ with nonzero mass, if it exists; otherwise, set $y_{1}^{*}(t)=1$. Then $y_{1}^{*}(t)<y_{0}(t)$, and that $\pi_{t}(\mathsf{d}x)$ is supported on $U_{t}\subset\{y_{0}(t)\}\cup[-1,y_{1}^{*}(t)]$. Therefore we have
$\int_{\mathbb{R}}\left(1+t+2\sqrt{t}x\right)^{m}\pi_{t}(\mathsf{d}x)=\\ \int_{\{y_{0}(t)\}}\left(1+t+2\sqrt{t}x\right)^{m}\pi_{t}(\mathsf{d}x)+\int_{-1}^{y_{1}^{*}(t)}\left(1+t+2\sqrt{t}x\right)^{m}\pi_{t}(\mathsf{d}x).$ (2.34)
The first term on the RHS equals $(1+t+2\sqrt{t}y_{0}(t))^{m}\pi_{t}(\{y_{0}(t)\})$. The absolute value of the second term on the RHS is bounded by $(1+t+2\sqrt{t}y_{1}^{*}(t))^{m}\left\lVert\pi_{t}\right\rVert_{TV}$, which has a lower order as $m\rightarrow\infty$ since $1<1+t+2\sqrt{t}y_{1}^{*}(t)<1+t+2\sqrt{t}y_{0}(t)$. Therefore by (2.33),
$\mathbb{E}_{\lambda_{m}}\left[t^{\tau_{1}+\dots+\tau_{m}}\right]=(t\rho+1-\rho)^{m}\sim\frac{A^{m}}{(1+A)^{2m}}\left(1+t+2\sqrt{t}y_{0}(t)\right)^{m}\pi_{t}(\{y_{0}(t)\}),$ (2.35)
where we write $f(m)\sim g(m)$ if $f(m)/g(m)\rightarrow 1$ as $m\rightarrow\infty$. The above implies
$t\rho+1-\rho=\frac{A}{(1+A)^{2}}\left(1+t+2\sqrt{t}y_{0}(t)\right)=\frac{1+At}{1+A}$
hence $\rho=A/(1+A)$, and also $\pi_{t}(\{y_{0}(t)\})=1$. Therefore, we conclude that (2.35) is an equality, and hence the second term on the RHS of (2.34) is equal to zero. If $y_{1}^{*}(t)>1$ is an atom, then by a similar analysis as above, this second term
$0=\int_{-1}^{y_{1}^{*}(t)}\left(1+t+2\sqrt{t}x\right)^{m}\pi_{t}(\mathsf{d}x)\sim\left(1+t+2\sqrt{t}y_{1}^{*}(t)\right)^{m}\pi_{t}(\{y_{1}^{*}(t)\}),$
which is a contradiction since $\pi_{t}(\{y_{1}^{*}(t)\})\neq 0$ by our assumption. Therefore we conclude that $\pi_{t}(\mathsf{d}x)$ only has a single atom at $y_{0}(t)$ with mass $1$. Observe that the continuous part density (2.2) of an Askey–Wilson signed measure has fixed sign over $x\in[-1,1]$, we conclude that the continuous part of $\pi_{t}(\mathsf{d}x)$ is constantly zero. Using formula (2.2) we have
$\left(q,t,AC,AD,C/A,D/A,CD/t\right)_{\infty}=0.$
Since $t\in(1-\varepsilon,1)$, $A>C\geq 0$ and $D\in(-1,0]$, we have $ACq^{k}=1$ for some $k\in\mathbb{N}_{0}$.
We first assume $k\in\mathbb{Z}_{+}$. We choose $t\in(1-\varepsilon,1)$ satisfying $A\sqrt{t}q\neq 1$. Then by $ACq^{k}=1$, we have either $A\sqrt{t}q>1$ or $C/\sqrt{t}>1$. When $A\sqrt{t}q>1$, by formula (2.4), the second largest atom generated by $A\sqrt{t}$, i.e., $\frac{1}{2}\left(A\sqrt{t}q+\frac{1}{A\sqrt{t}q}\right)$, has nonzero mass. When $C/\sqrt{t}>1$, by (2.3), the largest atom generated by $C/\sqrt{t}$, i.e., $\frac{1}{2}\left(\frac{C}{\sqrt{t}}+\frac{\sqrt{t}}{C}\right)$, has nonzero mass. This contradicts to the fact that $\pi_{t}(\mathsf{d}x)$ has only one atom $y_{0}(t)$ with nonzero mass.
We conclude that $k=0$, i.e., $AC=1$. When $AC=1$, by a computation, the multi-dimensional Askey–Wilson signed measure $\pi^{(A,1/A,C,D)}_{t_{1},\dots,t_{m}}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)$ on the RHS of (1.6) is a point mass at $x_{i}=y_{0}(t_{i})$ for $i=1,\dots,m$. Using (1.6) we conclude that $\lambda_{m}$ is the product Bernoulli measure on $\{0,1\}^{m}$ with density $A/(1+A)$. Therefore $\lambda$ is the product Bernoulli measure on $\{0,1\}^{\mathbb{Z}_{+}}$ with the same density. We conclude the proof. ∎
Appendix A Bounding the total variation distance by generating functions
In this appendix, we provide a general bound of the total variation distance between two probability measures on $\{0,1\}^{m}$ by the value of the difference of their joint generating functions at certain points. While it is possible that this inequality is already known, we have been unable to locate it in the previous literature.
Proposition A.1.
Let $\kappa$ and $\kappa^{\prime}$ be probability measures on $\{0,1\}^{m}$. Then for any set of numbers $0<t_{i,0}<t_{i,1}$ for $i=1,\dots,m$, we have:
$\mathsf{d}_{TV}(\kappa,\kappa^{\prime})\leq\frac{1}{2}\prod_{i=1}^{m}\frac{1+t_{i,1}}{t_{i,1}-t_{i,0}}\sum_{\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}}\left\lvert\mathbb{E}_{\kappa}\left[\prod_{i=1}^{m}t_{i,\upsilon_{i}}^{\tau_{i}}\right]-\mathbb{E}_{\kappa^{\prime}}\left[\prod_{i=1}^{m}t_{i,\upsilon_{i}}^{\tau_{i}}\right]\right\lvert,$ (A.1)
where $\tau_{i}\in\{0,1\}$ is the occupation variable on the site $i$, for $i=1,\dots,m$.
Proof.
This result can be seen as a simple corollary of the following inequality: For any degree-$1$ multivariate polynomial $G$ with real coefficients:
$G(x_{1},\dots,x_{m})=\sum_{\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}}g_{\upsilon_{1},\dots,\upsilon_{m}}x_{1}^{\upsilon_{1}}\dots x_{m}^{\upsilon_{m}},$
and for any set of numbers $0<t_{i,0}<t_{i,1}$, $i=1,\dots,m$, we have:
$\sum_{\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}}|g_{\upsilon_{1},\dots,\upsilon_{m}}|\leq\prod_{i=1}^{m}\frac{1+t_{i,1}}{t_{i,1}-t_{i,0}}\sum_{\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}}|G(t_{1,\upsilon_{1}},\dots,t_{m,\upsilon_{m}})|.$ (A.2)
Specifically, (A.1) can be seen by taking
$g_{\tau_{1},\dots,\tau_{m}}=\kappa(\tau_{1},\dots,\tau_{m})-\kappa^{\prime}(\tau_{1},\dots,\tau_{m})$
for any $\tau_{1},\dots,\tau_{m}\in\{0,1\}$.
We now prove (A.2) by induction on $m$. When $m=1$, we have $G(x_{1})=g_{0}+g_{1}x_{1}$, where
$g_{0}=\frac{-t_{1,0}G(t_{1,1})+t_{1,1}G(t_{1,0})}{t_{1,1}-t_{1,0}},\quad g_{1}=\frac{G(t_{1,1})-G(t_{1,0})}{t_{1,1}-t_{1,0}}.$
Therefore
$|g_{0}|+|g_{1}|\leq\frac{1+t_{1,1}}{t_{1,1}-t_{1,0}}\left(|G(t_{1,0})|+|G(t_{1,1})|\right).$
Suppose $m\geq 2$, and (A.2) holds for the $m-1$ case. We write
$G(x_{1},\dots,x_{m})=G_{0}(x_{1},\dots,x_{m-1})+G_{1}(x_{1},\dots,x_{m-1})x_{m},$
where for $j\in\{0,1\}$,
$G_{j}(x_{1},\dots,x_{m-1})=\sum_{\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}}g_{\upsilon_{1},\dots,\upsilon_{m-1},j}x_{1}^{\upsilon_{1}}\dots x_{m-1}^{\upsilon_{m-1}}.$
By the inequality for $m=1$, for any $0<t_{m,0}<t_{m,1}$ and real values of $x_{1},\dots,x_{m-1}$ we have:
$|G_{0}(x_{1},\dots,x_{m-1})|+|G_{1}(x_{1},\dots,x_{m-1})|\\ \leq\frac{1+t_{m,1}}{t_{m,1}-t_{m,0}}\left(|G(x_{1},\dots,x_{m-1},t_{m,0})|+|G(x_{1},\dots,x_{m-1},t_{m,1})|\right).$
We sum the above inequality over all the $2^{m-1}$ possibilities $x_{i}\in\{t_{i,0},t_{i,1}\}$, $11\leq i\leq m-$ 1 and get:
$\sum_{j\in\{0,1\}}\sum_{\upsilon_{1},\dots,\upsilon_{m-1}\in\{0,1\}}|G_{j}(t_{1,\upsilon_{1}},\dots t_{m-1,\upsilon_{m-1}})|\\ \leq\frac{1+t_{m,1}}{t_{m,1}-t_{m,0}}\sum_{\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}}|G(t_{1,\upsilon_{1}},\dots t_{m,\upsilon_{m}})|.$ (A.3)
By the induction hypothesis, for each $j\in\{0,1\}$,
$\sum_{\upsilon_{1},\dots,\upsilon_{m-1}\in\{0,1\}}|g_{\upsilon_{1},\dots,\upsilon_{m-1},j}|\leq\prod_{i=1}^{m-1}\frac{1+t_{i,1}}{t_{i,1}-t_{i,0}}\sum_{\upsilon_{1},\dots,\upsilon_{m-1}\in\{0,1\}}|G_{j}(t_{1,\upsilon_{1}},\dots t_{m-1,\upsilon_{m-1}})|.$
We sum the above inequality over $j\in\{0,1\}$ and get:
$\begin{split}\sum_{\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}}|g_{\upsilon_{1},\dots,\upsilon_{m}}|&\leq\prod_{i=1}^{m-1}\frac{1+t_{i,1}}{t_{i,1}-t_{i,0}}\sum_{j\in\{0,1\}}\sum_{\upsilon_{1},\dots,\upsilon_{m-1}\in\{0,1\}}|G_{j}(t_{1,\upsilon_{1}},\dots t_{m-1,\upsilon_{m-1}})|\\ &\leq\prod_{i=1}^{m}\frac{1+t_{i,1}}{t_{i,1}-t_{i,0}}\sum_{\upsilon_{1},\dots,\upsilon_{m}\in\{0,1\}}|G(t_{1,\upsilon_{1}},\dots t_{m,\upsilon_{m}})|,\end{split}$
where the last step uses (A.3). Hence inequality (A.2) is proved for $m$ case. We conclude the proof. ∎
Appendix B Total variation bounds of Askey–Wilson signed measures
In this appendix, we establish two results which bound the total variations of certain Askey–Wilson signed measures.
Proposition B.1.
Assume $A,C\geq 0$, $-1<B,D\leq 0$ and $q\in[0,1)$. Assume also that $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$ and that $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exist positive constants $K\geq 1$ and $\varepsilon$ depending on $A,B,C$ and $q$ but not on $D$, such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$, the total variation of the Askey–Wilson signed measure $P_{s,t}(x,\mathsf{d}y)$ is bounded from above by $\frac{K}{(t-s)^{2}}$.
Remark B.2.
As mentioned in Remark 2.14, the above total variation bound is more refined than the one utilized in [30]. In particular, the total variation bound of $P_{s,t}(x,\mathrm{d}y)$ provided by [30, Proposition A.1] can be derived as a simple corollary of the above result, but the converse does not hold. As we will see in the proof, more subtle estimates of the atom masses are needed to establish this result.
Before we offer the proof of Proposition B.1, we first prove a lemma bounding the total variation of an Askey–Wilson signed measure by the supremum of all the atom masses:
Lemma B.3.
Assume $q\in[0,1)$ and $(a,b,c,d)\in\Omega$, where $\Omega$ is introduced in Definition 2.1. Then
$\left\lVert\nu\left(\mathsf{d}y;a,b,c,d\right)\right\rVert_{TV}\leq 1+2\operatorname{card}\left(F(a,b,c,d)\right)\sup_{y_{j}^{\mathbf{e}}\in F(a,b,c,d)}|p_{j}^{\mathbf{e}}(a,b,c,d)|,$ (B.1)
where $\left\lVert\nu\right\rVert_{TV}$ denotes the total variation of a signed measure $\nu$ and $\operatorname{card}(\mathcal{S})$ denotes the cardinality of a set $\mathcal{S}$. We recall that $p_{j}^{\mathbf{e}}(a,b,c,d)$ denotes the mass of the atom $y_{j}^{\mathbf{e}}$.
Proof.
We recall from Section 2.1 that for any Askey–Wilson signed measure, the continuous part density on $(-1,1)$ must be either constantly positive or constantly negative. Furthermore, the total mass is always equal to $1$. Therefore we have,
$\left\lVert\nu\left(\mathsf{d}y;a,b,c,d\right)\right\rVert_{TV}\leq\sum_{y_{j}^{\mathbf{e}}\in F(a,b,c,d)}|p_{j}^{\mathbf{e}}|+\int_{-1}^{1}|f(y;a,b,c,d)|\mathsf{d}y\leq 1+2\sum_{y_{j}^{\mathbf{e}}\in F(a,b,c,d)}|p_{j}^{\mathbf{e}}|.$
The RHS above can be further bounded by the RHS of (B.1). ∎
Proof of Proposition B.1.
We assume that $A,C>0$ in the proof. Otherwise we are in the fan region $AC<1$, in which the Askey–Wilson signed measures are actually probability measures, whose total variation equal $1$. We first prove that there exists $K\geq 1$ and $\varepsilon>0$ depending on $A,B,C,D$ and $q$, such that for any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$,
$\left\lVert P_{s,t}(x,\mathsf{d}y)\right\rVert_{TV}\leq\frac{K}{(t-s)^{2}}.$ (B.2)
At the end of the proof we will show that the above constants $K$ and $\varepsilon$ can be chosen as independent of $D$.
We start by choosing $\varepsilon>0$ according to (the proof of) Theorem 2.5. For any $s<t$ in $I=[1,1+\varepsilon)$ and $x\in U_{s}$, we will investigate the Askey–Wilson signed measure $P_{s,t}^{(A,B)}(x,\mathsf{d}y)=\nu\left(\mathsf{d}y;a,b,c,d\right)$, where
$a=A\sqrt{t},\quad b=B\sqrt{t},\quad c=\sqrt{\frac{s}{t}}\left(x+\sqrt{x^{2}-1}\right),\quad d=\sqrt{\frac{s}{t}}\left(x-\sqrt{x^{2}-1}\right).$ (B.3)
Notice that the norms of $a,b,c,d$ are uniformly bounded by a finite constant. Hence the total number of atoms in $\nu\left(\mathsf{d}y;a,b,c,d\right)$ is also uniformly bounded. Here and below, a uniform constant means a constant that only depends on $A,B,C,D$ and $q$. In view of Lemma B.3 we only need to bound the supremum of norms of all atom masses by a uniform constant over $(t-s)^{2}$.
We look at the atom masses: If $|aq^{k}|\geq 1$ for $k\in\mathbb{N}_{0}$, then:
$\begin{split}p_{0}^{\mathbf{a}}&=\frac{(a^{-2},bc,bd,cd)_{\infty}}{(b/a,c/a,d/a,abcd)_{\infty}},\\ p_{k}^{\mathbf{a}}&=\frac{(a^{-2},bc,bd,cd)_{\infty}}{(b/a,c/a,d/a,abcd)_{\infty}}\frac{q^{k}(1-a^{2}q^{2k})(a^{2},ab,ac,ad)_{k}}{(q)_{k}(1-a^{2})a^{4k}\prod_{l=1}^{k}\left((b/a-q^{l})(c/a-q^{l})(d/a-q^{l})\right)},\quad k\geq 1.\end{split}$ (B.4)
The masses for atoms generated by $\mathbf{e}\in\{\mathbf{c},\mathbf{d}\}$ are given by similar formulas with $a$ and $\mathfrak{e}$ swapped. Note that $\mathbf{b}$ does not generate atoms since $b=B\sqrt{t}\in(-1,0]$. One can observe that the numerators of all atom masses are uniformly bounded from above. In the denominator, the term $(abcd)_{\infty}$ is uniformly bounded away from 00, since $abcd$ equals $ABt\leq 0$. Other three terms $(1-\mathfrak{e}^{2})$, $\mathfrak{e}^{4k}$ and $(q)_{k}$ are uniformly bounded away from 00, since we have $|\mathfrak{e}q^{k}|\geq 1$ and that $k\geq 1$ is bounded from above. Therefore, for any atom $y_{k}^{\mathbf{e}}$ generated by $\mathbf{e}\in\{\mathbf{a},\mathbf{c},\mathbf{d}\}$,
$|p_{k}^{\mathbf{e}}(a,b,c,d)|\leq\operatorname{Const}\times\prod_{\mathfrak{f}\in\{a,b,c,d\}\setminus\{\mathfrak{e}\}}\prod_{l=-k}^{\infty}\frac{1}{|1-q^{l}\mathfrak{f}/\mathfrak{e}|},\quad k\in\mathbb{N}_{0},\quad|\mathfrak{e}q^{k}|\geq 1.$ (B.5)
Here and below, we use $\operatorname{Const}$ to denote a uniform positive constant. The specific value of this constant may vary in different bounds.
The bound (B.2) then follows from the following claim:
Claim B.4.
For any $\mathbf{e}\in\{\mathbf{a},\mathbf{c},\mathbf{d}\}$ such that $|\mathfrak{e}|\geq 1$, we have $\mathfrak{e}\geq 1$. Consider the following set:
$\mathcal{A}_{k}^{\mathbf{e},\mathbf{f}}:=\left\{\left|1-q^{l}\frac{\mathfrak{f}}{\mathfrak{e}}\right|:l\geq-k\right\},\quad k\in\mathbb{N}_{0},\quad\mathfrak{e}q^{k}\geq 1,\quad\mathbf{f}\in\{\mathbf{a},\mathbf{b},\mathbf{c},\mathbf{d}\}\setminus\{\mathbf{e}\}.$
For fixed $\mathbf{e}$, $\mathbf{f}$ and $k$, all the elements in the set $\mathcal{A}_{k}^{\mathbf{e},\mathbf{f}}$ are uniformly bounded away from 00 except for possibly one ‘exceptional’ element. For this exceptional element $|1-q^{l}\mathfrak{f}/\mathfrak{e}|$, we have that either $\mathbf{f}=\mathbf{a}$ or $\mathbf{f}=\mathbf{c}$, and that it can be lower bounded by $(t-s)$ times a uniform positive constant.
Notice that, although the RHS of (B.5) is an infinite product, we do not need to worry about its tail, because it is uniformly bounded as the inverse of a $q$-Pochhammer symbol. If Claim B.4 holds, then for any fixed $\mathbf{e}\in\{\mathbf{a},\mathbf{c},\mathbf{d}\}$ and $k\in\mathbb{N}_{0}$, there are at most two pairs of $\mathbf{f}\in\{\mathbf{a},\mathbf{c}\}$ and $l\geq-k$ that produce exceptional elements $|1-q^{l}\mathfrak{f}/\mathfrak{e}|$. Hence in view of (B.5), the atom mass $|p_{k}^{\mathbf{e}}(a,b,c,d)|$ is bounded by a uniform constant times $1/(t-s)^{2}$. Therefore in view of Lemma B.3, bound (B.2) holds.
The proof of Claim B.4 necessitates a detailed analysis of all the possible values of $\mathfrak{f}/\mathfrak{e}$, which is provided in the following claim:
Claim B.5.
For any $\mathfrak{e}\in\{a,c,d\}$ satisfying $|\mathfrak{e}|\geq 1$, we have $\mathfrak{e}\geq 1$. Consider the value of $\mathfrak{f}/\mathfrak{e}$ for $\mathfrak{f}\in\{a,b,c,d\}\setminus\{\mathfrak{e}\}$ with $\mathfrak{f}>0$. Either $|\mathfrak{f}/\mathfrak{e}|$ can be expressed as an element in
$\left\{\frac{\sqrt{s}}{At},\frac{s}{t},\frac{t}{s},\frac{1}{A^{2}t},\frac{C}{At},\frac{At}{C}\right\}$
times $q^{r}$ for some $r\in\mathbb{Z}$, or $\mathfrak{e}=c$, $\mathfrak{f}=d$ and $\mathfrak{f}/\mathfrak{e}\leq q^{2k}s/t$, where $k\in\mathbb{N}_{0}$ satisfies $\mathfrak{e}q^{k}\geq 1$.
Proof of Claim B.5.
We recall that we always have $a>0$ and $b\in(-1,0]$. Similar to the proof of Theorem 2.5, we consider the following three cases depending on $x\in U_{s}$:
  1. Case 1.
    Let $x\in[-1,1]$ then we have $a=A\sqrt{t}$, $b=B\sqrt{t}$, $c=\sqrt{\frac{s}{t}}\left(x+\sqrt{x^{2}-1}\right)$ and $d=\sqrt{\frac{s}{t}}\left(x-\sqrt{x^{2}-1}\right)$. We have $|c|,|d|<1$ and $|c/a|=|d/a|=\sqrt{s}/(At)$. Claim B.5 holds.
  2. Case 2.
    Let $x=\frac{1}{2}\left(q^{j}A\sqrt{s}+\left(q^{j}A\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}A\sqrt{s}>1$. Then $a=A\sqrt{t}$, $b=B\sqrt{t}$, $c=q^{j}As/\sqrt{t}>0$ and $d=1/\left(q^{j}A\sqrt{t}\right)\in(0,1)$. We have $c/a=q^{j}s/t$, $d/a=1/(q^{j}A^{2}t)$, $a/c=t/(q^{j}s)$ and $d/c=1/(q^{2j}A^{2}s)$. If $\mathfrak{e}=c$ then $q^{k}c=q^{k+j}As/\sqrt{t}\geq 1$, and hence $d/c\leq q^{2k}s/t$. Therefore Claim B.5 holds.
  3. Case 3.
    Let $x=\frac{1}{2}\left(q^{j}C/\sqrt{s}+\left(q^{j}C/\sqrt{s}\right)^{-1}\right)$ for $j\in\mathbb{N}_{0}$ and $q^{j}C/\sqrt{s}>1$. Then $a=A\sqrt{t}$, $b=B\sqrt{t}$, $c=q^{j}C/\sqrt{t}>0$ and $d=s/(q^{j}C\sqrt{t})\in(0,1)$. We have $c/a=q^{j}C/(At)$, $d/a=s/(q^{j}ACt)$, $a/c=At/(q^{j}C)$ and $d/c=s/(q^{2j}C^{2})$. If $\mathfrak{e}=c$ then $q^{k}c=q^{k+j}C/\sqrt{t}\geq 1$, and hence $d/c\leq q^{2k}s/t$. Therefore Claim B.5 holds.
We conclude the proof of Claim B.5. ∎
We now begin to prove Claim B.4.
Proof of Claim B.4.
The fact that $\mathfrak{e}\geq 1$ has been proven by Claim B.5. To prove the rest of the statements, we use Claim B.5 to divide the proof into the following three cases:
  1. Case 1.
    When $\mathfrak{f}\leq 0$, we have $|1-q^{l}\mathfrak{f}/\mathfrak{e}|\geq 1$. In this case, every element in $\mathcal{A}_{k}^{\mathbf{e},\mathbf{f}}$ is lower bounded by $1$.
  2. Case 2.
    When we have
    $\frac{\mathfrak{f}}{\mathfrak{e}}\in\left\{\frac{\sqrt{s}}{At},\frac{s}{t},\frac{t}{s},\frac{1}{A^{2}t},\frac{C}{At},\frac{At}{C}\right\}\times\{q^{r}:r\in\mathbb{Z}\},$
    where we denote $A\times B:=\{ab:a\in A,b\in B\}$ for $A,B\subset\mathbb{R}$. We factorize the numbers above as products of the parts that involve $A$ and $C$ and the other parts that involve $s$ and $t$ (which can be $1$):
    $\frac{\sqrt{s}}{At}=\frac{1}{A}\times\frac{\sqrt{s}}{t},\quad\frac{s}{t}=1\times\frac{s}{t},\quad\frac{t}{s}=1\times\frac{t}{s},\quad\\ \frac{1}{A^{2}t}=\frac{1}{A^{2}}\times\frac{1}{t},\quad\frac{C}{At}=\frac{C}{A}\times\frac{1}{t},\quad\frac{At}{C}=\frac{A}{C}\times t.$
    One can shrink the value of $\varepsilon>0$ if necessary, to ensure that there exists an uniform $\kappa>0$ such that for each $x\in\{1/2,1,-1\}$ and $y=\pm 1$ and for any $1\leq s<t<1+\varepsilon$ we have $|1-q^{i}us^{x}t^{y}|>\kappa$ when: (1) $i\in\mathbb{Z}\setminus\{0\}$ and $u=1$ or (2) $i\in\mathbb{Z}$ and
    $u\in\left\{\frac{1}{A},\frac{1}{A^{2}},\frac{C}{A},\frac{A}{C}\right\}\setminus\{q^{j}:j\in\mathbb{Z}\}.$
    Therefore, in the cases when
    $\frac{\mathfrak{f}}{\mathfrak{e}}\in\left\{\frac{s}{t},\frac{t}{s}\right\}\times\{q^{r}:r\in\mathbb{Z}\}:$
    Denote $\frac{\mathfrak{f}}{\mathfrak{e}}\in\left\{q^{h}\frac{s}{t},q^{h}\frac{t}{s}\right\}$, $h\in\mathbb{Z}$. Then if $h+l\in\mathbb{Z}\setminus\{0\}$ we have $\left|1-q^{l}\frac{\mathfrak{f}}{\mathfrak{e}}\right|\geq\kappa$. If $h+l=0$,
    $\left|1-q^{l}\frac{\mathfrak{f}}{\mathfrak{e}}\right|\geq\min\left(\left|1-\frac{s}{t}\right|,\left|1-\frac{t}{s}\right|\right)\geq\operatorname{Const}\times(t-s).$
    By a similar argument, in the other cases when
    $\frac{\mathfrak{f}}{\mathfrak{e}}\in\left\{\frac{\sqrt{s}}{At},\frac{1}{A^{2}t},\frac{C}{At},\frac{At}{C}\right\}\times\{q^{r}:r\in\mathbb{Z}\},$
    then $\left|1-q^{l}\frac{\mathfrak{f}}{\mathfrak{e}}\right|$ can be lower bounded by $\operatorname{Const}\times(t-s)$ if
    $q^{-l}\in\left\{\frac{1}{A},\frac{1}{A^{2}},\frac{C}{A},\frac{A}{C}\right\}\quad\text{and}\quad\frac{\mathfrak{f}}{\mathfrak{e}}=\left\{\frac{\sqrt{s}}{t},\frac{1}{t},t\right\}\times q^{-l};$
    and for all other $l\in\mathbb{Z}$, $\left|1-q^{l}\frac{\mathfrak{f}}{\mathfrak{e}}\right|$ can be lower bounded by $\kappa$. The above fact follows from
    $\min\left(\left|1-\frac{\sqrt{s}}{t}\right|,\left|1-\frac{1}{t}\right|,|1-t|\right)\geq\operatorname{Const}\times(t-s),$
    and from our choices of $\varepsilon>0$ and $\kappa>0$.
  3. Case 3.
    When $\mathfrak{e}=c$, $\mathfrak{f}=d$ and $\mathfrak{f}/\mathfrak{e}\leq q^{2k}s/t$. In this case $q^{l}\mathfrak{f}/\mathfrak{e}\leq q^{2k+l}s/t<q^{2k+l}$. If either $k\geq 1$ or $k=0$ and $l\geq 1$, then $\mathfrak{f}/\mathfrak{e}<q$ and $|1-\mathfrak{f}/\mathfrak{e}|>1-q$. If $k=l=0$ then $\mathfrak{f}/\mathfrak{e}\leq s/t$ and
    $\left|1-\frac{\mathfrak{f}}{\mathfrak{e}}\right|\geq 1-\frac{s}{t}\geq\operatorname{Const}\times(t-s).$
In view of Case 1, Case 2 and Case 3 above, we conclude the proof of Claim B.4. ∎
We return to the proof. By the reasoning below Claim B.4, we conclude the proof of (B.2). It is clear from our choice of $\varepsilon>0$ that it is independent of $D$. Our last goal is to show that the constant $K$ can also be chosen as not depending on $D$. Since $D\in(-1,0]$, $\max(|a|,|b|,|c|,|d|)$ can be bounded by a constant not depending on $D$. Therefore the total number of atoms $\operatorname{card}\left(F(a,b,c,d)\right)$ as well as the constant prefactors appearing in the atom mass bounds (B.5) can be bounded by constants independent of $D$. In the proof of Claim B.4 (which bound all the linear factors $|1-q^{l}\mathfrak{f}/\mathfrak{e}|$), $D$ does not play a role at all. Therefore the uniform constants chosen from Claim B.4 does not depend on $D$. In summary, constant $K$ can be chosen as independent of $D$. We conclude the proof. ∎
Proposition B.6.
Assume $A,C\geq 0$, $\max(A,C)>1$, $-1<B,D\leq 0$, $q\in[0,1)$ and $|(ABCD)_{\infty}|\leq 1$. Assume also $A/C\notin\{q^{l}:l\in\mathbb{Z}\}$ if $A,C\geq 1$. Then there exists a positive constant $L$ depending on $A,C$ and $q$ such that
$\left\lVert\pi_{1}\right\rVert_{TV}\leq\frac{L}{1-BD}\left\lvert\pi_{1}\left(\{y_{0}(1)\}\right)\right\lvert.$
Proof.
We will prove the result for the high density phase $A>1$, $A>C$. The result for the low density phase follows from symmetry. The atoms are generated by $A$ and possibly also by $C$. In view of Lemma B.3 we want to bound the mass $\left\lvert\pi_{1}\left(\{y_{0}(1)\}\right)\right\lvert$ of first atom from below and masses of all other atoms from above. We have:
$\pi_{1}\left(\{y_{0}(1)\}\right)=p_{0}^{\mathbf{a}}(A,B,C,D)=\frac{(A^{-2},BC,BD,CD)_{\infty}}{(B/A,C/A,D/A,ABCD)_{\infty}}.$
Notice that on $x\in(-\infty,0]$, $(x)_{\infty}\geq 1$ is a decreasing function. Using the assumption $-1<B,D\leq 0$, one can lower bound this atom mass:
$|\pi_{1}\left(\{y_{0}(1)\}\right)|\geq P\frac{1-BD}{|(ABCD)_{\infty}|},$ (B.6)
where $P$ is a positive constant depending only on $A,C$ and $q$.
We then look at the masses of other (possible) atoms generated by $\mathbf{a}$: For $k\geq 1$ and $Aq^{k}\geq 1$,
$p_{k}^{\mathbf{a}}(A,B,C,D)\\ =\frac{(A^{-2},BC,BD,CD)_{\infty}}{(B/A,C/A,D/A,ABCD)_{\infty}}\frac{q^{k}(1-A^{2}q^{2k})(A^{2},AB,AC,AD)_{k}}{(q)_{k}(1-A^{2})A^{4k}\prod_{l=1}^{k}\left((B/A-q^{l})(C/A-q^{l})(D/A-q^{l})\right)},$
along with the masses of the (possible) atoms generated by $\mathbf{c}$: For $k\geq 0$ and $Cq^{k}\geq 1$,
$p_{k}^{\mathbf{c}}(A,B,C,D)\\ =\frac{(C^{-2},AB,BD,AD)_{\infty}}{(B/C,A/C,D/C,ABCD)_{\infty}}\frac{q^{k}(1-C^{2}q^{2k})(C^{2},BC,AC,CD)_{k}}{(q)_{k}(1-C^{2})C^{4k}\prod_{l=1}^{k}\left((B/C-q^{l})(A/C-q^{l})(D/C-q^{l})\right)}.$
Again, using the assumption $-1<B,D\leq 0$, one can upper bound these atom masses:
$|p_{k}^{\mathbf{e}}(A,B,C,D)|\leq\frac{Q}{|(ABCD)_{\infty}|},$ (B.7)
for any atom $y_{k}^{\mathbf{e}}\in F(A,B,C,D)$, where $Q$ is a uniform positive constant depending on $A,C$ and $q$.
In view of Lemma B.3, the lower bound (B.6), the upper bound (B.7) and our assumption that $|(ABCD)_{\infty}|\leq 1$, we conclude the proof.∎
Appendix C Time-reversal of Askey–Wilson signed measures
We prove a special symmetry of multi-dimensional Askey–Wilson signed measures known as the time-reversal. In the case that this symmetry only involves actual probability measures (i.e., on the fan region $AC<1$), this symmetry appears as [9, equation (5.10)]. However the proof is not explained clearly enough therein. Here we adopt a different approach to prove the result for the general case.
Proposition C.1.
Assume $q\in[0,1)$ and $A,B,C,D\in\mathbb{R}$ satisfying $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$. We have:
$\pi_{t_{1},\dots,t_{m}}^{(A,B,C,D)}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)=\pi_{1/t_{m},\dots,1/t_{1}}^{(C,D,A,B)}\left(\mathsf{d}x_{m},\dots,\mathsf{d}x_{1}\right)$ (C.1)
for any $0<t_{1}\leq\dots\leq t_{m}$ for which the multi-dimensional Askey–Wilson signed measures on both sides of this identity are well-defined.
Remark C.2.
In view of the definition (2.7) of $\pi_{t_{1},\dots,t_{m}}^{(A,B,C,D)}\left(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m}\right)$, it is well-defined when $\pi^{(A,B,C,D)}_{t_{1}}(\mathsf{d}y)$ and each $P^{(A,B)}_{t_{i-1},t_{i}}(x,\mathsf{d}y)$, $i=2,\dots,m$ for any $x\in U_{t_{i-1}}$ are well-defined Askey–Wilson signed measures, i.e. their quadruples of entries lie in the subset $\Omega\subset\mathbb{C}^{4}$ in Definition 2.1.
Before we begin the proof of Proposition C.1, we recall the orthogonality and the projection formula of Askey–Wilson polynomials and signed measures. The Askey–Wilson polynomials $w_{j}(x):=w_{j}(x;a,b,c,d)$, $j\in\mathbb{N}_{0}$ are defined by three term recurrence in [1] (see also [30, Section 2.1]).
Lemma C.3 (Corollary 2.6 and Theorem 2.7 in [30]).
Assume $a,b,c,d\in\Omega$. Then for any $j,k\in\mathbb{N}_{0}$,
$\int_{\mathbb{R}}\nu(\mathsf{d}x;a,b,c,d)w_{j}(x)w_{k}(x)=\delta_{jk}\frac{(1-q^{j-1}abcd)(q,ab,ac,ad,bc,bd,cd)_{j}}{(1-q^{2j-1}abcd)(abcd)_{j}}.$ (C.2)
Lemma C.4 (Proposition 3.3 in [30]).
Assume $A,B,C,D\in\mathbb{R}$. For any $s\leq t$ such that $P_{s,t}(x,\mathsf{d}y)$ is well-defined for all $x\in U_{s}$, we have:
$\int_{\mathbb{R}}p_{j}(y;t)P_{s,t}(x,\mathsf{d}y)=p_{j}(x;s),$ (C.3)
where for $j\in\mathbb{N}_{0}$,
$p_{j}(x;t):=t^{j/2}(ABt)_{j}^{-1}w_{j}\left(x;A\sqrt{t},B\sqrt{t},C/\sqrt{t},D/\sqrt{t}\right).$
Remark C.5.
We note that Lemma C.3 and Lemma C.4 above are originally due to [1, 8] and later generalized by [30] to the cases of signed measures.
We now start to prove Proposition C.1.
Proof of Proposition C.1.
We first prove the result for $m=1$ and $m=2$, then we use induction to prove it for $m\geq 3$.
For $m=1$, we recall that the Askey–Wilson signed measures are symmetric with respect to their entries. This fact in particular follows from the Askey–Wilson polynomials $w_{j}(x)=w_{j}(x;a,b,c,d)$ being symmetric with respect to parameters $a,b,c$ and $d$. Therefore we have:
$\pi_{t}^{(A,B,C,D)}(\mathsf{d}x)=\nu\left(\mathsf{d}x;A\sqrt{t},B\sqrt{t},\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}}\right)\\ =\nu\left(\mathsf{d}x;\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}},A\sqrt{t},B\sqrt{t}\right)=\pi_{1/t}^{(C,D,A,B)}(\mathsf{d}x).$
For $m=2$, we compute, for $s<t$ and $j,k\in\mathbb{N}_{0}$:
$\begin{split}&\int_{\mathbb{R}^{2}}\pi_{s,t}^{(A,B,C,D)}\left(\mathsf{d}x,\mathsf{d}y\right)w_{j}\left(y;A\sqrt{t},B\sqrt{t},\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}}\right)w_{k}\left(x;A\sqrt{s},B\sqrt{s},\frac{C}{\sqrt{s}},\frac{D}{\sqrt{s}}\right)\\ &=\frac{(ABt)_{j}(ABs)_{k}}{t^{j/2}s^{k/2}}\int_{\mathbb{R}^{2}}\pi_{s,t}^{(A,B,C,D)}\left(\mathsf{d}x,\mathsf{d}y\right)p_{j}(y;t)p_{k}(x;s)\\ &=\frac{(ABt)_{j}(ABs)_{k}}{t^{j/2}s^{k/2}}\int_{\mathbb{R}}\pi_{s}^{(A,B,C,D)}(\mathsf{d}x)p_{k}(x;s)\int_{\mathbb{R}}P^{(A,B)}_{s,t}(x,\mathsf{d}y)p_{j}(y;t)\\ &=\frac{(ABt)_{j}(ABs)_{k}}{t^{j/2}s^{k/2}}\int_{\mathbb{R}}\pi_{s}^{(A,B,C,D)}(\mathsf{d}x)p_{k}(x;s)p_{j}(x;s)\\ &=\frac{(ABt)_{j}s^{j/2}}{(ABs)_{j}t^{j/2}}\int_{\mathbb{R}}\pi_{s}^{(A,B,C,D)}(\mathsf{d}x)w_{j}\left(x;A\sqrt{s},B\sqrt{s},\frac{C}{\sqrt{s}},\frac{D}{\sqrt{s}}\right)w_{k}\left(x;A\sqrt{s},B\sqrt{s},\frac{C}{\sqrt{s}},\frac{D}{\sqrt{s}}\right)\\ &=\frac{(ABt)_{j}s^{j/2}}{(ABs)_{j}t^{j/2}}\delta_{jk}\frac{(1-q^{j-1}ABCD)(q,ABs,AC,AD,BC,BD,CD/s)_{j}}{(1-q^{2j-1}ABCD)(ABCD)_{j}}\\ &=\delta_{jk}\left(\frac{s}{t}\right)^{j/2}(ABt)_{j}\left(\frac{CD}{s}\right)_{j}\frac{(1-q^{j-1}ABCD)(q,AC,AD,BC,BD)_{j}}{(1-q^{2j-1}ABCD)(ABCD)_{j}}.\end{split}$ (C.4)
We have used the projection formula (Lemma C.4) in the third step and the orthogonality property (Lemma C.3) in the second to last step.
The RHS of the above identity (C.4) remains the same when we swap $x\leftrightarrow y$, $j\leftrightarrow k$, $A\leftrightarrow C$, $B\leftrightarrow D$ and also take $s\mapsto 1/t$, $t\mapsto 1/s$. Therefore we have:
$\begin{split}&\int_{\mathbb{R}^{2}}\pi_{s,t}^{(A,B,C,D)}\left(\mathsf{d}x,\mathsf{d}y\right)w_{j}\left(y;A\sqrt{t},B\sqrt{t},\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}}\right)w_{k}\left(x;A\sqrt{s},B\sqrt{s},\frac{C}{\sqrt{s}},\frac{D}{\sqrt{s}}\right)\\ &=\delta_{jk}\left(\frac{s}{t}\right)^{j/2}(ABt)_{j}\left(\frac{CD}{s}\right)_{j}\frac{(1-q^{j-1}ABCD)(q,AC,AD,BC,BD)_{j}}{(1-q^{2j-1}ABCD)(ABCD)_{j}}\\ &=\int_{\mathbb{R}^{2}}\pi_{1/t,1/s}^{(C,D,A,B)}\left(\mathsf{d}y,\mathsf{d}x\right)w_{k}\left(x;\frac{C}{\sqrt{s}},\frac{D}{\sqrt{s}},A\sqrt{s},B\sqrt{s}\right)w_{j}\left(y;\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}},A\sqrt{t},B\sqrt{t}\right)\\ &=\int_{\mathbb{R}^{2}}\pi_{1/t,1/s}^{(C,D,A,B)}\left(\mathsf{d}y,\mathsf{d}x\right)w_{j}\left(y;A\sqrt{t},B\sqrt{t},\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}}\right)w_{k}\left(x;A\sqrt{s},B\sqrt{s},\frac{C}{\sqrt{s}},\frac{D}{\sqrt{s}}\right).\end{split}$
We notice that $w_{j}(x)$ has degree $j$ for $j\in\mathbb{N}_{0}$. One can use induction within the above identity to conclude that, for every $j,k\in\mathbb{N}_{0}$,
$\int_{\mathbb{R}^{2}}\pi_{s,t}^{(A,B,C,D)}\left(\mathsf{d}x,\mathsf{d}y\right)x^{k}y^{j}=\int_{\mathbb{R}^{2}}\pi_{1/t,1/s}^{(C,D,A,B)}\left(\mathsf{d}y,\mathsf{d}x\right)x^{k}y^{j}.$
We next choose a large enough closed ball $K\subset\mathbb{R}^{2}$ whose interior contains the supports of both of the signed measures $\pi_{s,t}^{(A,B,C,D)}\left(\mathsf{d}x,\mathsf{d}y\right)$ and $\pi_{1/t,1/s}^{(C,D,A,B)}\left(\mathsf{d}y,\mathsf{d}x\right)$. By the Stone–Weierstrass theorem, any continuous function $g$ on $K$ can be uniformly approximated by polynomials. Therefore for any continuous function $g$ on $\mathbb{R}^{2}$, we have:
$\int_{\mathbb{R}^{2}}\pi_{s,t}^{(A,B,C,D)}\left(\mathsf{d}x,\mathsf{d}y\right)g(x,y)=\int_{\mathbb{R}^{2}}\pi_{1/t,1/s}^{(C,D,A,B)}\left(\mathsf{d}y,\mathsf{d}x\right)g(x,y).$ (C.5)
Hence both of the signed measures define the same bounded linear functional
$C(K)\longrightarrow\mathbb{R}.$
In view of (the uniqueness part of) the Riesz representation theorem (see for example [26, Theorem 6.19]), we have
$\pi_{s,t}^{(A,B,C,D)}\left(\mathsf{d}x,\mathsf{d}y\right)=\pi_{1/t,1/s}^{(C,D,A,B)}\left(\mathsf{d}y,\mathsf{d}x\right).$
We conclude the proof for $m=2$.
Assume that the result holds for some $m\geq 2$, we prove it for the case $m+1$. We have:
$\begin{split}&\pi^{(A,B,C,D)}_{t_{1},\dots,t_{m+1}}(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m+1})\\ &=\pi^{(A,B,C,D)}_{t_{1},\dots,t_{m}}(\mathsf{d}x_{1},\dots,\mathsf{d}x_{m})P^{(A,B)}_{t_{m},t_{m+1}}(x_{m},\mathsf{d}x_{m+1})\\ &=\pi^{(C,D,A,B)}_{1/t_{m},\dots,1/t_{1}}(\mathsf{d}x_{m},\dots,\mathsf{d}x_{1})P^{(A,B)}_{t_{m},t_{m+1}}(x_{m},\mathsf{d}x_{m+1})\\ &=\pi^{(C,D,A,B)}_{1/t_{m}}(\mathsf{d}x_{m})P^{(C,D)}_{1/t_{m},1/t_{m-1}}(x_{m},\mathsf{d}x_{m-1})\dots P^{(C,D)}_{1/t_{2},1/t_{1}}(x_{2},\mathsf{d}x_{1})P^{(A,B)}_{t_{m},t_{m+1}}(x_{m},\mathsf{d}x_{m+1})\\ &=\pi^{(A,B,C,D)}_{t_{m}}(\mathsf{d}x_{m})P^{(A,B)}_{t_{m},t_{m+1}}(x_{m},\mathsf{d}x_{m+1})P^{(C,D)}_{1/t_{m},1/t_{m-1}}(x_{m},\mathsf{d}x_{m-1})\dots P^{(C,D)}_{1/t_{2},1/t_{1}}(x_{2},\mathsf{d}x_{1})\\ &=\pi^{(A,B,C,D)}_{t_{m},t_{m+1}}(\mathsf{d}x_{m},\mathsf{d}x_{m+1})P^{(C,D)}_{1/t_{m},1/t_{m-1}}(x_{m},\mathsf{d}x_{m-1})\dots P^{(C,D)}_{1/t_{2},1/t_{1}}(x_{2},\mathsf{d}x_{1})\\ &=\pi^{(C,D,A,B)}_{1/t_{m+1},1/t_{m}}(\mathsf{d}x_{m+1},\mathsf{d}x_{m})P^{(C,D)}_{1/t_{m},1/t_{m-1}}(x_{m},\mathsf{d}x_{m-1})\dots P^{(C,D)}_{1/t_{2},1/t_{1}}(x_{2},\mathsf{d}x_{1})\\ &=\pi^{(C,D,A,B)}_{1/t_{m+1},\dots,1/t_{1}}(\mathsf{d}x_{m+1},\dots,\mathsf{d}x_{1}),\end{split}$
where we have used the result for $m$, $1$ and $2$. By induction, we conclude the proof. ∎